### v1

# Amputation-imputation based generation of synthetic tabular data for ratemaking

Abstract
Actuarial ratemaking depends on high-quality data, yet access to such data is often limited by the cost of obtaining new data, privacy concerns, etc. In this paper, we explore synthetic-data generation as a potential solution to these issues. In addition to discussing generative methods previously studied in the actuarial literature, we introduce to the insurance community another approach based on Multiple Imputation by Chained Equations (MICE). We present a comparative study using an open-source dataset and evaluating MICE-based models against other generative models like Variational Autoencoders and Conditional Tabular Generative Adversarial Networks. We assess how well synthetic data preserves the original marginal distributions of variables as well as the multivariate relationships among covariates. We also investigate the consistency between Generalized Linear Models (GLMs) trained on synthetic data with GLMs trained on the original data. Furthermore, we assess the ease of use of each generative approach and study the impact of augmenting original data with synthetic data on the performance of GLMs for predicting claim counts. Our results highlight the potential of MICE-based methods in creating high-quality tabular data while being more user-friendly than the other methods.
Keywords: synthetic actuarial data, data augmentation, multiple imputation by chained equations, generative adversarial networks, autoencoders
1 Introduction
Motivation. High-quality actuarial data is essential for actuaries in various business lines of insurance and reinsurance companies. It helps quantify risks more accurately, and thus maintain the competitiveness of actuarial pricing and reserving models. Furthermore, researchers in actuarial science need realistic data to test their novel methodologies and compare them with existing ones. Even if insurers and reinsurers may have enough data for some products or business lines, the companies rarely publicly share their data due to concerns related to data privacy, security, and their competitive advantage. As a result, the number of open-access actuarial datasets is quite small (see, e.g., Gabrielli and Wüthrich, (2018), Cote et al., (2020), Segura-Gisbert et al., (2024)). Furthermore, the companies themselves may also face a lack of data, e.g., when they enter a new market or release a new (re-)insurance product.
To increase the amount of actuarial data, one can generate synthetic data, which can be used as a standalone dataset or as an enlargement to the existing dataset. In the former case, synthetic data can be publicly shared and used for research purposes (e.g., testing new models and algorithms), if it has the statistical properties of the real data and does not expose sensitive information. As for the latter case, the augmentation of real data by synthetic data allows to train models on a larger and potentially more diverse dataset and, as a result, improve their capacity to generalize to unseen data. It can be especially helpful for actuaries when collecting and processing additional real data is impossible or very costly in terms of time or money.
There are two types of approaches to generating synthetic data: those that have assumptions regarding the distribution of data to be generated and those that do not impose such assumptions. We will focus on the latter type, as one of our aims is to develop an actuarial data generator that can be easily used “out-of-the-box”. Furthermore, the insurance application we target is general ratemaking (without telematics), where, usually, datasets are tabular with both numeric and categorical variables and temporal independence of observations is assumed.
The generation of synthetic data without the need for distributional assumptions gained much popularity due to the advancement of deep generative models such as generative adversarial networks (GANs) and variational autoencoders (VAEs). To our knowledge, only three papers investigated the use of such models in the context of ratemaking: Kuo, (2020), Cote et al., (2020), Jamotton and Hainaut, (2024). We discuss these articles in more detail in the literature overview below and mention for now only aspects that strongly motivate us to conduct further research on this topic. First, in our experience, the data generation approaches studied in those three papers seem to require quite some customization for each input dataset, which is why their ease of use in daily work of pricing actuaries seems to be limited. Second, the benefit of augmenting original data with synthetic data is not addressed in those papers.
Our contributions. First, we introduce to actuarial science several amputation-imputation-based approaches to data generation. These techniques are based on Multiple Imputation by Chained-Equations (MICE) with Random Forests (RFs) as imputation models. According to our tests, these approaches can be more suitable for “out-of-the-box” usage in practice than the generative models studied in Kuo, (2020), Cote et al., (2020), Jamotton and Hainaut, (2024), while having a comparable performance.
Second, using a publicly available French Motor Third-Part Liability dataset called freMTPL2freq (Dutang and Charpentier,, 2019) as a training dataset, we compare $9$ approaches to ratemaking data generation, including the existing approaches (Kuo,, 2020; Cote et al.,, 2020; Jamotton and Hainaut,, 2024), their modifications that address the challenge of generative adversarial networks (GANs) in dealing with high-cardinality categorical variables, and the newly introduced amputation-imputation-based methods. We analyze the quality of synthetic datasets by studying two aspects: the similarity between the distribution of the original data and of synthetic data, the ability of methods to capture structural dependence among covariates and the response variable. Furthermore, we assess the impact of augmenting original data with synthetic data on the performance of GLMs for predicting claim frequency.
Our main results are threefold. First, we illustrate that the MICE-RF approach is a competitive method for generation of synthetic ratemaking data in comparison to deep generative models such as a VAE and a CTGAN. The important practical advantage of the MICE-RF algorithm is its ease of use, given the existing packages in $\mathrm{R}$ and the versatility of RFs in dealing with variables of different types without much pre-processing. Second, we show that modifying a CTGAN with an AE for categorical-variable pre-processing improves the quality of synthetic high-cardinality categorical variables but deteriorates the performance of this hybrid approach in terms of other criteria. In general, a MICE-RF approach and a CTGAN without AE seem to be the most balanced in terms of the performance metrics we consider. Third, our case studies indicate that for our experimental data sets based on freMTPL2freq data augmentation does not improve the performance of GLMs trained on the original data merged with synthetic one.
Literature overview. Kuo, (2020) pioneered the use of conditional tabular GANs for generating actuarial datasets, focusing on ratemaking and lapse modeling. In his experiments with freMTPL2freq dataset, the researcher showed that generalized linear models (GLMs) trained on synthetic data achieved performance comparable to GLMs trained on real data, and the distributions of both continuous and categorical variables in the synthetic datasets closely mirrored the real datasets. However, the relativities (how much each level of a categorical variable affects the predicted outcome compared to a reference level) of GLM models constructed using synthetic datasets do not show the same patterns as those of models created using real datasets. Because of the observed instability in the parameters of a GLM trained on synthetic data, developing a rating plan based on it might not be the best approach.
Inspired by Kuo, (2020), Cote et al., (2020) investigated three GAN-based models for generating synthetic insurance data: Multi-Categorical Wasserstein GAN with Gradient Penalty (MC-WGAN-GP), Conditional Tabular GAN (CTGAN), and Mixed Numerical and Categorical Differentially Private GAN (MNCDP-GAN). In a case study using the freMTPLfreq data, they find that MC-WGAN-GP performs best in replicating univariate distributions and the observed claim frequency for each category of categorical variables, followed by CTGAN, and then MNCDP-GAN. In assessing the consistency of GLMs fitted on real versus synthetic data, CTGAN-based GLMs most closely resemble the real-data-based GLM in terms of average estimated GLM coefficients. However, MC-WGAN-GP based GLM predictions are closest to the real-data-based GLM predictions when using Mean Absolute Error and Mean Squared Error as evaluation metrics. The authors conclude that neither of the tested models performs well across all key performance indicators (KPIs) with a high level of training data confidentiality.
As an alternative to GAN-based models, Jamotton and Hainaut, (2024) propose a Variational Autoencoder (VAE) model for the generation of insurance data. Their approach features a quantile transformation for continuous variables to address multimodality. In their VAE, the researchers use a reconstruction loss function that consists of two parts: cross-entropy for categorical variables and MSE for continuous variables. In a case study using a data set from the homepage of the book Ohlsson and Johansson, (2010), the researchers find that the marginal distributions of the real datasets and those of the synthetic dataset are almost the same, as supported by visual inspection, Kolmogorov-Smirnov tests for continuous variables, and Chi-Squared tests for categorical variables. Based on Pearson and Spearman correlations for continuous variables and contingency scores based on contingency tables for categorical variables, the authors conclude that synthetic data preserves pairwise relationships among variables reasonably well. Training a GLM on synthetic data generated by their VAE and noticing that the respective Poisson deviance is almost the same as the one for a GLM trained on the original data, the scientists conclude that synthetic data can substitute real data in actuarial analyses.
Fundamentally different data generation methods are those based on data imputation and have not yet been explored in the actuarial science context. The main idea behind those methods is to sequentially “amputate” parts of data and then impute it with a suitable statistical learning model. To our knowledge, this idea was first introduced by D.B. Rubin, who first developed a data imputation approach called MICE in the 1970s (cf. Rubin, (1977)) then proposed to use this approach also to create (fully or partially) synthetic datasets in Rubin, (1993). We refer readers interested in the subsequent development of MICE-based generation of synthetic data to Drechsler, (2011) and emphasize two aspects. First, this approach gained popularity in various scientific domains (e.g., see Volker and Vink, (2021) for medicine), yet it seems to have been overseen by the actuarial community. Second, the imputation in this framework is done on a variable-by-variable (column-by-column) basis, and the imputation model can be flexibly chosen from linear regression to sophisticated models.
A related idea of generating artificial data using imputation appeared in Neves et al., (2022), where the authors propose a so-called tabulator framework. In their approach, they propose to randomly (not on variable-by-variable basis) select data entries to be amputated and then imputed by a so-called back-end engine for imputation. As a model driving this imputation engine the authors propose to use one of the three methods that can be seen as modifications of Generative Adversarial Imputation Networks (GAINs, originally proposed by Yoon et al., (2018)): Slim GAIN (SGAIN), Wasserstein Slim GAIN with Clipping Penalty (WSGAIN-CP), Wasserstein Slim GAIN with Gradient Penalty (WSGAIN-GP). According to the researchers, tabulator has a comparable performance (tested on 10 datasets from various fields like healthcare, economics, and computer science) with CTGAN that, in turn, is better than medGAN, VEEGAN, and table-GAN. The researchers conclude that the tabulator preserves the dependence structure of the original dataset better than CTGAN, but both approaches struggle when generating synthetic data from original imbalanced data. It is also stated that synthetic data from tabulator improves the performance of regression/classification tasks only marginally.
Readers interested in the literature on data generation beyond traditional ratemaking are referred to So et al., (2021) and So and Jeong, (2022) for telematics data, to Gabrielli and Wüthrich, (2018), Avanzi et al., (2021), Avanzi et al., (2023), Bauwelinckx et al., (2024) for reserving data, to Campo and Antonio, (2023) for auto-insurance fraud-detection data, and to Chapter 10 in Wuthrich et al., (2025) for a brief introduction to generative modeling tools for actuaries. We refer readers to Fonseca and Bacao, (2023) and Brophy et al., (2023) for a systematic review of the literature on the GAN-based generation of tabular data and on the GAN-based generation of time series data, respectively. For an overview comparing deep generative models, including but not limited to VAEs and GANs, we direct readers to Bond-Taylor et al., (2022).
Structure of the paper. In Section 2, we present the general setup for tabular data generation in the context of ratemaking. In Section 3, we describe the main components of data generation approaches compared in this paper. Section 4 is devoted to our comparative study. Section 5 concludes the paper. Appendix A contains additional information on the hybrid approach combining CTGAN with AEs.
2 General setup
First, we introduce notation and briefly discuss some data pre-processing techniques commonly used before fitting/training any statistical learning models. Second, we summarize the basics of GLMs for predicting insurance claim frequencies. Third, we explain what is meant by generating synthetic data and state two research questions that are answered in the subsequent sections.
We denote a tabular dataset by $\mathcal{S}:=\left\{\bm{s}_{i}\right\}_{i=1,\dots,n_{ROD}}$, where $n_{ROD}$ is the number of rows in the original data (number of observations), $\bm{s}_{i}=(\bm{x}_{i},v_{i},y_{i})$ is a tuple with three elements such that $\bm{x}_{i}=(x_{i;1},\dots,x_{i;j},\dots,x_{i;n_{C}})^{\top}$ is a $n_{C}$-dimensional vector of observed values of covariates (independent variables), $n_{C}\in\mathbb{N}$, is the transpose operator, $v_{i}\in(0,1]$ is an exposure-to-risk (proportion of a year during which the observation $i$ was tracked) and $y_{i}$ is an observed value of an one-dimensional response variable (dependent variable) for anobservation $i=1,\dots,n_{ROD}$. A tuple $\bm{s}_{i}$ can be interpreted as an insurance policy snippet, $\bm{x}_{i}$ – as a vector of policy characteristics, $v_{i}$ – as the fraction of a year during which a policy snippet $i$ was active, and $y_{i}$ – as the observed number of claims (or the monetary loss associated with a claim in the case of severity modeling). For simplicity, we may omit the index $i$ indicating the ordinal number of the observation, when it is not essential for interpretation, and just use $(\bm{x},v,y)$. The vector $\bm{x}$ can have both categorical and numerical components. We denote by $n_{CC}$ the number of categorical covariates in $\bm{x}$ and by $n_{NC}=n_{C}-n_{CC}$ the number of numerical covariates in $\bm{x}$.
Before fitting/training any model on the raw data, it is usually pre-processed to a suitable format, depending on the model to be used later. After pre-processing, a variable can be represented by multiple elements, which are referred to as features. An example of pre-processing of a categorical variable is one-hot encoding. Consider a categorical variable $x_{j}$, $j=1,\dots,n_{CC}$, which has $m_{j}$ different categories (labels) $\left\{a_{1}^{j},\dots,a_{m_{j}}^{j}\right\}$. The one-hot encoding of this variable means representing it as an $m_{j}$-dimensional vector containing only 0 and $1$ as follows:
$x_{j}\mapsto\bm{x}_{j}^{cat}=\left(x_{j_{1}},\dots,x_{j_{m_{j}}}\right)^{\top}=\left(\mathbbm{1}_{\left\{x_{j}=a_{1}^{j}\right\}},\dots,\mathbbm{1}_{\left\{x_{j}=a_{m_{j}}^{j}\right\}}\right)^{\top}\in\left\{0,1\right\}^{m_{j}},$ (2.1)
where $\mathbbm{1}_{\left\{\cdot\right\}}$ is an indicator function that is equal to $1$ if the condition in curly brackets is true and 0 otherwise. One-hot encoding is often used (implicitly or explicitly) before training advanced statistical learning models, including those mentioned in the next section. Another way of encoding categorical variables is dummy encoding, which yields a similar representation as (2.1), but one of the categories is omitted. This omitted category (label) is then called a reference one and is represented by an $(m_{j}-1)$-dimensional vector containing only 0. Dummy encoding is frequently used for GLMs and GAMs.
Unless specified otherwise, we assume that for data generation algorithms, categorical variables are pre-processed via one-hot encoding and for GLMs – via dummy encoding. We denote by $n_{PCF}$ the number of pre-processed categorical features.
Similarly, statistical learning models tend to perform better when numerical variables are also pre-processed, which should be explicitly done by the user before training a model or is implicitly done in a respective package implementing the model. For example, a numerical variable $x_{k}$ (as well as exposure $v$ and response variable $y$) can be transformed via a min-max scaler to $\widetilde{x}_{k}\in[-1,1]$, $k=1,\dots,n_{NC}=:n_{PNF}$, where $PNF$ stands for pre-processed numerical features.
Thus, without loss of generality (w.l.o.g.) and with a slight abuse of notation, we can write the vector of pre-processed features as follows:
$\bm{x}=\left((\bm{x}^{cat})^{\top},(\bm{x}^{num})^{\top}\right)^{\top}:=\left((\bm{x}_{1}^{cat})^{\top},\dots,(\bm{x}_{n_{CC}}^{cat})^{\top},\widetilde{x}_{1},\dots,\widetilde{x}_{n_{PNF}}\right)^{\top}\in\mathbb{R}^{n_{PF}},$
where $n_{PF}=n_{PCF}+n_{PNF}$ stands for the number of pre-processed features.
To avoid loaded notation, we still use $\mathcal{S}$ to denote the dataset after pre-processing. An actuary can use a subset of $\mathcal{S}_{train}\subseteq\mathcal{S}$ (where subsetting is done among tuples not within their distinct elements) to train a generative model that, once trained, can be used to generate a synthetic dataset $\mathcal{S}^{A}=\left\{{\bm{s}}^{A}_{i}\right\}_{i=1,\dots,n_{RSD}}$, where the superscript $A$ refers to the approach chosen for generative modeling, $n_{RSD}$ is the number of needed rows of synthetic data and ${\bm{s}}^{A}_{i}=\left({\bm{x}}^{A}_{i},{v}_{i}^{A},{y}_{i}^{A}\right)$. A good approach $A$ should ensure that the distribution $F_{\bm{s}}$ of a real (pre-processed) data tuple is close to the distribution $F_{{\bm{s}}^{A}}$ of a synthetic (pre-processed) data tuple. This implies that the data generation approach should learn the structural link between $(\bm{x},v)$ and $y$.
In ratemaking, it is common to model the link between $y$ and $(\bm{x},v)$ via a generalized linear model (GLM). GLMs were introduced in Nelder and Wedderburn, (1972) and became very popular in insurance, where they are routinely used in various business lines such as claim management, underwriting, etc. In our paper, we will only discuss the basics of a Poisson GLM, which is frequently used to predict insurance claim counts and which we use in numerical studies. For a deep dive into the theory and practice of GLMs in insurance readers are referred to Chapter 4 of Denuit and Trufin, (2019) and Chapter 5 of Wüthrich and Merz, (2023).
Let $Y$ be a random variable representing a response variable (e.g., number of insurance claims per year) and $y$ be its realization (observation). Let $\bm{X}$ be a vector of (random) features in a format suitable for a GLM (e.g., categorical variables are represented via dummy encoding, the first feature always equals $1$ if a GLM shall contain an intercept, etc.) and $\bm{x}$ be its realization (observation). Let $v\in(0,1]$ denote a known exposure. Then a Poisson GLM with a so-called log-link function and without interactions among covariates assumes that:
$Y|\bm{X}=\bm{x},v]\sim\text{Poisson}\left(v\cdot\exp\left(\bm{\beta}^{\top}\bm{x}\right)\right),$
where $\bm{\beta}:=\left(\beta_{0},\dots,\beta_{n_{PF}}\right)^{\top}$ is the vector of GLM parameters. Taking an expectation and applying a logarithmic function, we get:
$\ln\Bigl{(}\underbrace{\mathbb{E}[Y|\bm{X}=\bm{x},v]}_{=y,\text{ target}}\Bigr{)}=\underbrace{\ln(v)}_{\text{offset}}+\underbrace{\bm{\beta}^{\top}\bm{x}}_{\text{linear predictor}},$ (2.2)
where $\mathbb{E}[\cdot|\cdot]$ denotes the conditional expectation operator.
A Poisson GLM specified by (2.2) has no interactions. Depending on the data, it may be beneficial to include interactions in a GLM. For simplicity, let us consider a pairwise interaction. An interaction between two features $x_{j}$ and $x_{k}$ (elements of $\bm{x}$) is modeled as an additional term $I\left(x_{j},x_{k}\right)$ in the right-hand side of (2.2) such that $I\left(x_{j},x_{k}\right)$ cannot be represented as the sum of two univariate functions, i.e., $I\left(x_{j},x_{k}\right)\neq I_{1}(x_{j})+I_{2}(x_{k})$, $j\in\left\{1,\dots,n_{PF}\right\},k\in\left\{1,\dots,n_{PF}\right\}$, $j\neq k$. A common choice is
$I\left(x_{j},x_{k}\right)=\beta_{j,k}x_{j}x_{k}.$ (2.3)
In (2.3), $x_{j}$ (and/or $x_{k}$) can be just a feature representing one category. To consider an interaction with an entire categorical variable, a sum of different (2.3) (one for each category in the respective categorical variable, except for the dummy one) must be added to the right-hand side of (2.2). For more information on modeling interactions in GLMs, see, e.g., Fox and Weisberg, (2018), or Section 4.2.5 in Denuit and Trufin, (2019).
Let us denote by $\bm{\beta}^{\ast}$ the true value of $\bm{\beta}$. In practice, $\bm{\beta}^{\ast}$ is not known and must be estimated from collected data. Using a training data $\mathcal{S}_{train}$, an actuary can obtain an estimate of $\bm{\beta}^{\ast}$ by a so-called maximum likelihood estimation (MLE), which is equivalent to minimizing Poisson deviance in the case of a Poisson GLM. When comparing different GLM models, the one with a lower Poisson deviance on the respective test set $\mathcal{S}_{test}:=\mathcal{S}\setminus\mathcal{S}_{train}$ is considered a better-performing one. An alternative performance measure is the Akaike Information Criterion (AIC), which takes into account the number of parameters of a GLM and is based only on $\mathcal{S}_{train}$.
The predictive performance of a GLM depends on various aspects including but not limited to the choice of covariates used in a GLM and the quality as well as quantity of the observed data. Addressing the former aspect, an actuary should manually select explanatory variables (elements of $\bm{x}$) or obtain the list of important variables for a GLM using so-called variable-selection algorithms, such as backward, forward, or mixed variable selection (VS) methods. Let us denote by $\widehat{\bm{\beta}}:=\widehat{\bm{\beta}}\left(\mathcal{S}_{train}\right)$ the estimate of $\bm{\beta}^{\ast}$ when an actuary correctly specifies all variables with significant effects in a GLM, i.e., when GLM is trained with all elements of $\bm{x}$ that correspond to non-zero elements of $\bm{\beta}^{\ast}$. Let $\widehat{\widehat{\bm{\beta}}\,}:=\widehat{\widehat{\bm{\beta}}\,}\!\left(\mathcal{S}_{train}\right)$ denote the estimate of $\bm{\beta}^{\ast}$ when an actuary uses some VS method.
To address the aspect of data quantity and quality, an actuary can generate a synthetic dataset $\mathcal{S}^{A}$ of a desirable size and reasonable quality. In this situation, two natural questions arise: one is related to the usage of synthetic data as a standalone set, and the other one is related to working with the original dataset that is augmented with a part of the synthetic dataset. First, how well does $\mathcal{S}^{A}$ preserve (from training data $\mathcal{S}_{train}$) the marginal distributions of variables as well as the multivariate relationships among covariates and how consistent is a GLM trained on $\mathcal{S}^{A}$ with a GLM fitted on $\mathcal{S}_{train}$? Second, which advantages and disadvantages does training GLMs on an augmented dataset $\mathcal{S}_{train}\cup\mathcal{S}^{A}$ have in terms of GLM’s predictive performance and the estimates of $\bm{\beta}^{\ast}$ (e.g., $\widehat{\bm{\beta}}(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ and/or $\widehat{\widehat{\bm{\beta}}\,}\!(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ are closer to the true $\bm{\beta}^{\ast}$)? These are the two main research questions that we aim to answer in this paper.
3 Methodology
In this section, we give an overview of the main building blocks of different synthetic data generators explored in this paper. First, we start with the imputation framework MICE and explain how it can be used for synthetic data generation. Second, we briefly explain a CTGAN model. Third, we summarize deterministic and variational autoencoders (AEs).
3.1 Multiple imputation by chained equations
According to van Buuren, (2018), multiple imputation by chained equations (MICE) is one of the best approaches to dealing with missing data. MICE is also known as fully conditional specification, sequential regression multivariate imputation, or simply chained equations.
MICE commonly operates under the assumption that data is missing at random, which means that the probability of a value being missing depends only on observed values and not on unobserved values. Avoiding explicit assumptions on the joint distribution of the data, MICE works by specifying a set of conditional distributions for each variable with missing data, given the other variables in the data. The algorithm then iteratively imputes the missing values by drawing from these conditional distributions. This process of iterative drawing can be seen as a Gibbs sampler that will converge to the draws from the theoretical joint distribution of data, should this joint distribution exist (cf. Raghunathan et al., (2001)). In practice, the assumption about the existence of the joint distribution is often implicitly made.
Borrowing notation from Drechsler, (2011), we summarize the MICE algorithm:
  1. 1.
    For every missing value in a dataset, impute it using a simple technique and treat these initial imputed values as “placeholders”.
  2. 2.
    For a variable $Y_{j}$ set its “placeholders” back to missing.
  3. 3.
    Regress the observed values of $Y_{j}$ on other variables using some imputation model.
  4. 4.
    Replace missing values of $Y_{j}$ with predictions from the imputation model.
  5. 5.
    Repeat Steps 2-4 for each $j$ with missing data to complete one cycle.
  6. 6.
    Iterate Steps 2-4 for several cycles to update imputations until the convergence or the maximal number $n_{MI}$ of MICE iterations is reached.
  7. 7.
    Repeat Steps 1-6 $n_{IS}$ times (number of imputed sets) by changing the way inherent randomness is present in Steps 1 and 3.
Figure 1 visually illustrates the original MICE approach.
[…]
As for Step 7, the examples of factors contributing to the variability of the imputed datasets are: initial imputation of the missing data in Step 1 (mean, median, random draw from observed values, etc.), specification of the imputation model in Step 3 (e.g., hyperparameters, bootstrap sampling if using random forests, etc.).
Once multiple imputed datasets have been created, inference can be done regarding a quantity of interest (e.g., mean of a variable, correlation coefficient between two variables). So-called Rubin’s rules provide statistical foundation for such inference. For a scheme that illustrates how MICE works, see Figure 1. For more information, see, e.g., Section 5.1 in Drechsler, (2011).
As mentioned earlier, the MICE framework can be used to generate synthetic datasets. For simplicity, we assume that the training data has no missing values, otherwise, a 2-step approach is needed, where before the generation (in the second step) in the first step, missing values are imputed. The idea is to iteratively select parts of the original data and then replace them with synthetic values drawn from the posterior distribution learned by MICE.
There are different ways to sequentially choose chunks of data for replacement (imputation). One way is to choose a full column in each iteration, cf. Volker and Vink, (2021). This case has two peculiarities. First, the imputation model is trained on the observed data, which is then replaced by the model prediction. Second, the MICE algorithm converges immediately and each draw will be a direct draw from the posterior distribution, because a so-called monotone missingness pattern is satisfied for column-wise replacement/imputation (see Section 3.1.2 in Drechsler, (2011)). Another approach, which in general does not satisfy the monotone missingness pattern, is to randomly select in each amputation-imputation step cells whose values must be replaced. The values in these cells are deleted, interpreted as missing, and the MICE algorithm is executed, requiring more than one iteration to converge.
In our case studies, we test several ways of iteratively amputating data and then applying MICE to impute it. One of them is inspired by the tabulator approach of Neves et al., (2022), where the authors suggest to amputate data in chunks of approximately the same size and to impute deleted data using a GAIN. Instead of a GAIN, we will use MICE as the imputation engine.
Similarly to the MICE-based imputation, there are Rubin rules for analyzing (e.g., computing a point estimate, an estimate for the variance) unknown statistical quantities from fully synthetic datasets, cf. Reiter, (2003) and Section 6.1 in Drechsler, (2011).
In our case studies, we use random forests (RFs) as the imputation model. This imputation model and many others are implemented in the R-package mice, cf. van Buuren and Groothuis-Oudshoorn, (2011). RFs were introduced in Breiman, (2001). It is a non-parametric model that can capture complex relationships among variables without the need to specify a particular regression model. This makes RFs well-suited for imputing missing data in datasets with non-linear interactions among variables. Another important advantage of RFs is that they can often be used out-of-the-box and can be used rather easily to handle heterogeneous data. For more information on RFs in actuarial context, see, e.g., Chapter 4 in Denuit and Trufin, (2020) or Chapter 6 in Wuthrich et al., (2025).
3.2 Conditional Tabular GAN
In this subsection, which is based on Kuo, (2020) and Sections $2$ and $3$ of Cote et al., (2020), we describe how a GAN works and explain how a CTGAN extends it.
A GAN consists of two neural networks (NNs): a generator $G$ and a discriminator $D$. The generator creates synthetic data samples, while the discriminator evaluates their authenticity. Through this adversarial process, the generator learns to produce increasingly realistic synthetic data that can fool the discriminator.
Let us denote by $\bm{\theta}_{G}$ and $\bm{\theta}_{D}$ the vectors of parameters of $G$ and $D$ respectively. $G$ takes as input random noise $\bm{Z}\sim F_{\bm{Z}}$, where $F_{\bm{Z}}$ is its cumulative distribution function, and maps the input to the space of the data to be generated. Usually, $F_{\bm{Z}}$ is a multivariate standard normal distribution and the dimension of $\bm{Z}$ is lower than that of the data point to be generated. The generator’s output is a fake data point $G(\bm{Z};\bm{\theta}_{G})$, whose distribution we denote by $F_{G}$. The discriminator’s output $D(\bm{S};\bm{\theta}_{D})$ is a score between 0 and $1$, which can be interpreted as the probability of the input data point $\bm{S}$ to be a real data point.
A GAN-training process (search for optimal values of parameters $\theta_{G}$ and $\theta_{D}$) is a min-max game in the following sense. On the one hand, the discriminator is trained by maximizing the expected score $\mathbb{E}_{\bm{S}}\left[D(\bm{S};\bm{\theta}_{D})\right]$ of a real data point as well as minimizing (w.r.t. $\bm{\theta}_{D}$) the expected score of a synthetic data point $\mathbb{E}_{\bm{Z}}\left[D(G(\bm{Z};\bm{\theta}_{G});\bm{\theta}_{D})\right]$. On the other hand, the generator is trained by maximizing the discriminator’s expected score $\mathbb{E}_{\bm{Z}}\left[D(G(\bm{Z};\bm{\theta}_{G});\bm{\theta}_{D})\right]$ on a synthetic data point $\bm{Z}$. So the objective function of a GAN combines both goals and is given by:
$\min_{\bm{\theta}_{G}}\max_{\bm{\theta}_{D}}\left(\mathbb{E}_{\bm{S}}\left[\ln\left(D(\bm{S};\bm{\theta}_{D})\right)\right]+\mathbb{E}_{\bm{Z}}\left[\ln\left(1-D(G(\bm{Z};\bm{\theta}_{G});\bm{\theta}_{D})\right)\right]\right).$ (3.1)
To improve the convergence of optimization routines for (3.1), researchers have proposed modifications to (3.1) such as using Wasserstein-$1$ distance between the true distribution $F_{\bm{X}}$ and the synthetic distribution $F_{G}$, enforcing additional constraints, adding penalties, etc.
Proposed in Xu et al., (2019), CTGAN builds upon the GAN framework by incorporating several innovations to address the challenges of generating synthetic tabular data: mode-specific normalization, conditional generation of data, and handling of mixed data. Next we summarize an exemplary architecture of CTGAN, as in Cote et al., (2020).
As per Kuo, (2020), a CTGAN generates individual records by randomly selecting a variable (e.g., fuel type) and then choosing a value (e.g., diesel) based on the actual data frequency. The algorithm finds a matching training data row for this value and generates the remaining variables conditioned on the initial selection. Both the generated and true records are evaluated by the discriminator. Figure 3 in Cote et al., (2020) schematically illustrates this process and Figure 2 in Cote et al., (2020) visually exemplifies the architecture of the generator and discriminator. The discriminator (blue) and the generator (orange) each have two fully-connected layers to capture relationships between columns. The generator, which supports categorical variables, employs skip connections; see Xu et al., (2019) for extensions to continuous variables. CTGAN also uses the PacGAN framework in the discriminator to mitigate a so-called mode collapse. This peculiarity means that the model focuses on a small subset of the real data and, thus, has a bias towards the generation of samples that are not sufficiently diverse.
Although a CTGAN tends to be substantially better than a GAN for generating high-cardinality categorical variables, this type of variables can be still challenging for a CTGAN, depending on the characteristics of the real data such as the degree of cardinality and the frequency of each category. Aiming to improve the performance of a CTGAN for high-cardinality categorical variables, we try encoding them via non-linear autoencoders (AEs) before training a CTGAN.
3.3 Autoencoders
The original aim of AEs is learning a compressed representation of data (encoding) and then reconstructing the original input (decoding). Structurally, an AE can be thought of as a pair of connected feedforward NNs. So principal components analysis approach can be interpreted as a linear AE, i.e., the activation functions of both NNs are linear. Hinton and Salakhutdinov, (2006) demonstrated the power of non-linear AEs for learning better, lower-dimensional representations than linear AEs. In our paper, we use two different types of AEs due to their different application – deterministic AEs and variational AEs.
Our application of deterministic AEs is the transformation of high-cardinality categorical variables to low-dimensional numerical vectors (instead of the one-hot encoding) before fitting them to GAN-based models for data generation. The benefits of using such AEs for training NNs in insurance predictive modeling has been studied in Delong and Kozak, (2023), which is our main reference for this application of non-linear AEs.
Consider a vector of features $\bm{a}\in\mathbb{R}^{p}$, $p\in\mathbb{N}$. An AE is a representation learning model that consists of two functions – a so-called encoder $\varphi:\mathbb{R}^{p}\rightarrow\mathbb{R}^{l}$ and a decoder $\psi:\mathbb{R}^{l}\to\mathbb{R}^{p}$, $l\in\mathbb{N}$.
The mapping $\bm{a}\mapsto\varphi(\bm{a})$ yields a $l$-dimensional representation of the $p$-dimensional vector. The mapping $\overline{\bm{a}}\mapsto\psi(\overline{\bm{a}})$ tries to reconstruct the $p$-dimensional vector $\bm{a}$ from its $l$-dimensional representation $\overline{\bm{a}}=\varphi(\bm{a})$. The so-called reconstruction function of an AE is defined as
$\pi=\psi\circ\varphi:\mathbb{R}^{p}\to\mathbb{R}^{p}.$
Let $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be a function that measures the error (difference) between $\bm{a}$ and $\pi(\bm{a})$. Then for a training dataset $(\bm{a}_{i})_{i=1}^{n}$, the objective of an AE is to find the functions $\varphi$ and $\psi$ that minimize the so-called reconstruction error defined as follows:
$\frac{1}{n}\sum_{i=1}^{n}L(\pi(\bm{a}_{i}),\bm{a}_{i}).$ (3.2)
An AE is trained in an unsupervised manner. If an AE achieves a small reconstruction error, it is said that the encoder $\varphi$ extracts the most important information from a multi-dimensional vector of features $\bm{a}$. In our application, $\bm{a}$ will be a vector of features coming from a one-hot encoded categorical covariate, i.e., $\bm{a}=\bm{x}_{j}^{cat}$ with $\bm{x}_{j}^{cat}$ defined as per (2.1), $j\in\left\{1,\dots,n_{CC}\right\}$. For each categorical variable $\bm{x}_{j}^{cat}$, a separate AE is trained to reduce the dimensionality $p_{j}=m_{j}$ of $\bm{x}_{j}^{cat}$ to $l_{j}<p_{j}$, $j\in\left\{1,\dots,n_{CC}\right\}$. Afterwards, this representation $\overline{\bm{x}}_{j}^{cat}=\varphi\left({\bm{x}}_{j}^{cat}\right)$, instead of ${\bm{x}}_{j}^{cat}$, is used as input for CTGAN to generate data. The approach of using a separate AE for each categorical variable is called Separate AEs in Delong and Kozak, (2023). We use the same architecture of separate AEs, as proposed by Delong and Kozak, (2023). In particular, the reconstruction function $\pi$ is an NN with one hidden layer. The output-layer neurons have softmax activation functions and return the probabilities that the reconstructed feature takes a particular category. The category with the highest predicted probability is the one predicted for the reconstructed feature. $L$ in (3.2) is a cross-entropy loss function.
The approach described above is applied to all categorical features in the dataset. An example of the architecture of a neural network used in this paper to build the autoencoder of type Separate AEs for categorical features (with 2 and 3 categories) is presented in Figure 2 in Delong and Kozak, (2023). See their Section 3.2 for further details on Separate AEs.
Instead of using a separate AE for each categorical variable, we could have one joint AE for all categorical variables simultaneously. This approach is called Joint AE in Delong and Kozak, (2023). According to Figure 4 there, for a large number of epochs (e.g., $100$ for freMTPL2freq data) and a sufficiently high dimension of the representation of all categorical features (e.g., $15$ for freMTPL2freq data), the reconstruction error power measured with the cosine similarity is very similar for Separate AEs and Joint AE.
The second type of AEs used in our paper is variational autoencoders (VAEs). They were introduced by Kingma and Welling, (2013) and can be seen as a Bayesian extension to deterministic AEs. VAEs learn a probabilistic representation of data points in a latent space, which is why these models can be used for synthetic data generation in addition to representation learning. In our case studies, we also consider the VAE-based approach proposed by Jamotton and Hainaut, (2024), who also tested their methodology on a data set with about $62000$ observations. For an overview of VAEs and details about their methodology, we refer the reader to their article.
4 Case studies
This section is dedicated to our comparative study, which is based on freMTPL2freq dataset. In Subsection 4.1, we provide the details about our method training pipelines, the mechanics of data augmentation, and the evaluation pipelines for generation methods. In Subsection 4.2, we specify all the methods tested and state explicit formulas for modeling the structural dependence between observed covariates and simulated claim counts that represent “ground truth” in tested scenarios. The results are discussed in Subsection 4.3.
4.1 General setting and evaluation metrics
Our aim with this case study is to answer the research questions stated at the end of Section 2 and, in doing so, utilize pre-existing implementations of synthetic data generation methods with minimal user interventions for data preparation or method training. This aspect is significant in practice since, in our experience from the industry, actuaries seldom have sufficient time to implement and fine-tune their specific algorithms and methods. Since we want to identify best “off-the-shelf” data generation approach, we do not do any pre-processing of variables when preparing data in this section. Any processing is done by the generation method implementation. We execute our case study on the high-performance computation cluster at the University of Tartu and can provide our code upon request.
In our experiments, we use freMTPL2freq dataset to give us a baseline of what a realistic dataset might look like. It has $678013$ observations with $9$ explanatory variables ($5$ categorical and $4$ numeric), $1$ exposure variable and $1$ response variable (the number of claims). This dataset has been used in multiple actuarial data science articles, including Kuo, (2020), where it is used for training a CTGAN to generate synthetic data.
Although we have access to the original values of the observed response variable $y_{i}$ in the freMTPL2freq dataset, we do not know the true relation between the observed value $y_{i}$ and the corresponding observed covariates $\bm{x}_{i}$. To test whether the generation methods learn the relation correctly, we simulate $Y_{i}\sim Poisson\left(\exp\left(f(\bm{x}_{i})\right)\right)$, where $f(\bm{x}_{i})$ is a known function of covariates, and use the realized value as $y_{i}$. Thus, we implicitly assume that all observations have the same exposure $1$. We deem this assumption acceptable, since, in practice, the rate is set using the values for annual policies (e.g., yearly predicted number of claims), which are scaled to fit the duration desired by the client.
We test two specific forms of f{f}. First, we assume that f{f} is linear in all covariates $x_{j}$, $j\in\{0,\dots,d\}$, where $d+1$ is the number of terms (including the intercept) in the model. This yields ${f}_{lin}(\bm{x})=\displaystyle\sum_{j=0}^{d}\beta_{j}^{*}x_{j}$, where coefficients $\beta_{j}^{*}\in\mathbb{R}$ are known and serve as the “true” values defining the relationship between the response variable and the covariates in pre-processed form (cf. Section 2). From now on, ${f}_{lin}$ is referred to as the linear formula.
In the second formula we use the previously defined ${f}_{lin}$ and additionally assume a set of interactions among covariates, namely ${f}_{int}(\bm{x})={f}_{lin}(\bm{x})+\sum_{\{j,k\}\in\mathcal{I}}\beta_{j;k}x_{j}x_{k}$, where the interaction part is defined as in Equation (2.3) and $\mathcal{I}$ is the set of index pairs for the interactions of interest. Although we use double indices for notational simplicity, one can construct a simple mapping from two indices to single indices. The values ${f}_{lin}(\bm{x}_{i})$ and ${f}_{int}(\bm{x}_{i})$ are then used as the means of the respective Poisson distributions to simulate the number of claims $y_{i}$, $i\in\left\{1,\dots,n_{ROD}\right\}$.
So, in our experiments, we have two datasets, each of which contains simulated number claims for which we know their “true” structural dependence on original covariates. In the sequel, we refer to this artificially generated number of claims as the response variable and use the $90\%-10\%$ training-test split, i.e., the number of rows in the train set is $n_{RTR}=\lfloor n_{ROD}\cdot 0.9\rfloor$.
Let now $\mathcal{A}=\left\{A_{1},\dots,A_{n_{A}}\right\}$ be the set of algorithms (approaches) for generation of synthetic data, $n_{A}\in\mathbb{N}$ is the number of algorithms for synthetic data generation. For each approach $A$, we denote by $GM_{A}$ as the generation method for this algorithm.
For each algorithm, although the technical specifics of training can vary, the underlying pipeline we use is the same. We supply each algorithm the training data $\mathcal{S}_{train}$ with the artificially generated response variable and without altering any of the covariate variable values. Then each of the methods utilizes the supplied training data in accordance with its implementation and, thus, trains the generation method $GM_{A}$. However, for some methods AEs (see Section 3.3) are utilized to transform categorical variables. This requires additional steps in the training pipeline, which are outlined in Appendix A.
Each generation method is used to generate $n_{E}$ (referred to as the number of experiments) synthetic datasets, where the number of rows for a single synthetic dataset was the same as the training data. Denote a generated synthetic dataset as $\mathcal{S}^{A;k}$, where $k\in\{1,\dots,n_{E}\}$ and $A\in\mathcal{A}$.
Using the described pipeline, synthetic datasets $\mathcal{S}^{A;k}$ are generated for each algorithm $A\in\mathcal{A}$ and each $k\in\{1,\dots,n_{E}\}$. Then, for each synthetic dataset, we compute two types of metrics: the evaluation of the dataset metrics and the evaluation of model metrics.
With dataset metrics, we want to assess whether the proportion of data in some subset of synthetic data is the same or close to the proportion of data in the original training data. Statistical tests like Kolmogorov-Smirnov or Anderson-Darling for numeric variables and $\chi^{2}$-test for categorical variables could be used. However, if the data size is large, then these tests are very sensitive to the slightest differences between synthetic data and original training data. Therefore, we use other numeric metrics to see how well the proportions of data are captured across all variables and pair-wise combinations of variables.
For dataset evaluation, we compare the proportions of specific values in the original and synthetic data. The comparison of data is done on discretized forms of the original variables.
Denote the original variable $j$ before preprocessing by $x_{j}^{orig}$. If $x_{j}^{orig}$ is a categorical variable, then no additional steps need to be taken, since after preprocessing, the variable will be discretized into a vector of binary variables $\bm{x}_{j}^{cat}$.
If $x_{j}^{orig}$ is numeric, then it is discretized into $n_{bin_{j}}$ bins, where bin cut-off values are selected based on the quantiles of the numeric variable values in the original training data. The quantiles are used to ensure that the proportion of observations in each bin is the same or very close. Note that in the case when one value appears in several quantiles (e.g., one value is very prevalent in the data), then it has a bin all on its own. For each bin, then, a binary variable (feature) is added, and the vector of these features is denoted by $\bm{x}_{j}^{cat}$. Additionally, we denote by $\mathcal{I}_{j}$ the set of indices for preprocessed variables corresponding to $\bm{x}_{j}^{cat}$ .
For each feature $x_{k}$ from $\bm{x}_{j}^{cat}$ in training data, we calculate the ratio of observations belonging to that bin or category as $r\left(x_{k}\right)=\frac{\sum_{i=1}^{n_{RTR}}x_{i;k}}{n_{RTR}}$, where $k\in\mathcal{I}_{j}$. Similarly, we obtain $\widehat{r}\left(x_{u}\right)$ for any synthetic data.
These ratios are used to calculate two metrics for the original variable $x_{j}^{orig}$:
$MAE(x_{j}^{orig})=\frac{1}{\left|\mathcal{I}_{j}\right|}\displaystyle\sum_{k\in\mathcal{I}_{j}}\left|r\left(x_{k}\right)-\widehat{r}\left(x_{k}\right)\right|\quad\text{and}\quad MAPE(x_{j}^{orig})=\frac{1}{\left|\mathcal{I}_{j}\right|}\displaystyle\sum_{k\in\mathcal{I}_{j}}\frac{\left|r\left(x_{k}\right)-\widehat{r}\left(x_{k}\right)\right|}{\left|\widehat{r}\left(x_{k}\right)\right|},$
where $\left|\mathcal{I}_{j}\right|$ denotes the number of elements in set $\mathcal{I}_{j}$.
For pairwise data proportion assessment, we extend the above metrics to pairs of variables. Given two (original) variables $x_{a}^{orig}$ and $x_{b}^{orig}$, the ratio for any two features $x_{k}$ and $x_{u}$ with $k\in\mathcal{I}_{a}$ and $u\in\mathcal{I}_{b}$ in the training data is computed as $r\left(x_{k},x_{u}\right)=\frac{\sum_{i=1}^{n_{RTR}}x_{i;k}\cdot x_{i;u}}{n_{RTR}}$. For synthetic data, $\widehat{r}\left(x_{k},x_{u}\right)$ can be obtained in the same way. The metrics in this case are defined analogously to the single-variable case, except the summation is done over both sets of indices $\mathcal{I}_{a}$ and $\mathcal{I}_{b}$.
Lastly, for numerical variables, it is possible to assess if the correlations among them are retained by the data generation method. We calculate the pairwise correlations $\rho(x_{k},x_{u})$ for all numeric variable pairs $(x_{k},x_{u})$ in the training data. For synthetic data, we obtain the correlations in a similar way, and then calculate $MAE$ and $MAPE$ for correlations analogously to ratios above.
We use the synthetic datasets $\mathcal{S}^{A;k}$ and the training data $\mathcal{S}_{train}$ as the reference for dataset metrics evaluation. For model metrics, the scope is expanded to better understand what effect data augmentation has on predictive model performance.
Under the term data augmentation, we mean the creation of $m$ new datasets for each synthetic dataset $\mathcal{S}^{A;k}$. More precisely, given a synthetic dataset $\mathcal{S}^{A;k}$, we split it (row-wise) into $m$ disjoint (approximately) equally sized parts $\mathcal{S}_{l}^{A;k}$, $l\in\{1,\dots,m\},$ and append the increasing collection of these synthetic data parts to the original training data. In order to describe all possible combinations of training and synthetic data, we introduce a new structure parameter $\bm{s}=(t,L)$, where $t$ is the indicator stating whether the training set is included ($t=1$) or not ($t=0$), and $L$ is the number of included synthetic data parts. With this setup, we then get
$\mathcal{S}^{A;k;(0,L)}=\bigcup_{l=1}^{L}\mathcal{S}_{l}^{A;k}\text{ and }\mathcal{S}^{A;k;(1,L)}=\mathcal{S}_{train}\bigcup\mathcal{S}^{A;k;(0,L)},$
where $L\in\{1,\dots,m\}$, $\mathcal{S}^{A;k}=\bigcup_{l=1}^{m}\mathcal{S}_{l}^{A;k}$ and $\mathcal{S}_{j}^{A;k}\bigcap\mathcal{S}_{l}^{A;k}=\emptyset$, $\forall j\neq l$. By construction we have $\mathcal{S}^{A;k;(1,0)}=\mathcal{S}_{train}$ and $\mathcal{S}^{A;k;(0,m)}=\mathcal{S}^{A;k}$.
For the evaluation of model metrics, the underlying models need to be trained. We focus on using GLMs as the predictive model type. For each dataset $\mathcal{S}^{A;k;\bm{s}}$, two predictive GLM models are derived. One model, where the formula used to build the model is the same as the corresponding true model formula (i.e. linear formula or interactions formula), and one which is derived using a bidirectional step-wise variable selection using $AIC$. Note that for interaction formula, the scope for the variable selection included variables explicitly defining the interactions between the different variables (since this was the easiest way to include interactions in the step-wise search). The values related to the trained GLM models are denoted as follows:
  • $\bm{\beta}^{\ast}=\left(\beta_{0}^{\ast},\beta_{1}^{\ast},\dots,\beta_{d}^{\ast}\right)^{\top}$ – the vector of true model coefficients as defined in the respective response variable generation formula (i.e., linear formula or interaction formula).
  • $\widehat{\bm{\beta}}=\left(\widehat{\beta_{0}},\widehat{\beta_{1}},\dots,\widehat{\beta}_{d}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained on the training dataset $\mathcal{S}_{train}$.
  • $\widehat{\bm{\beta}}^{A;k}=\left(\widehat{\beta}_{0}^{A;k},\widehat{\beta}_{1}^{A;k},\dots,\widehat{\beta}_{d}^{A;k}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained only on the synthetic dataset $\mathcal{S}^{A;k}$.
  • $\widehat{\bm{\beta}}^{A;k;\bm{s}}=\left(\widehat{\beta}_{0}^{A;k;\bm{s}},\widehat{\beta}_{1}^{A;k;\bm{s}},\dots,\widehat{\beta}_{d}^{A;k;\bm{s}}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained on trained on the dataset specified by the data structure parameter $s$.
  • $MSE\left(\widehat{\beta}_{j}^{\bullet}\right)$ – mean squared error of the estimated coefficient $\widehat{\beta}_{j}^{\bullet}$, where $\bullet$” represents any algorithm/data structure combination.
As these values can be obtained for every trained GLM model, we use two metrics to assess the effect of synthetic data on estimated values of model coefficients.
First, to measure the distance between the true GLM coefficients and their estimate based on the true structure of a GLM and a synthetic dataset generated by an approach $A$ and specified by the structure parameter $\bm{s}$, we use the following metric:
$M_{1}\left(A,\bm{s}\right)=\frac{1}{d}\frac{1}{n_{E}}\sum_{k=1}^{n_{E}}\sum_{j=0}^{d}\frac{\left(\beta_{j}^{\ast}-\widehat{\beta}^{A;k;\bm{s}}_{j}\right)^{2}}{MSE\left(\widehat{\beta}_{j}\right)}.$ (4.1)
In (4.1), the denominator of each summand contains the mean squared error $MSE\left(\widehat{\beta}_{j}\right)$} for the true model coefficient estimated on the training data. This ensures that the distance values can be compared with each other.
The second metric we use is defined as follows:
$M_{2}\left(A,\bm{s}\right)=\frac{1}{d}\sum_{j=1}^{d}\frac{MSE\left(\widehat{\beta}_{j}\right)+\frac{1}{n_{E}}\sum_{k=1}^{n_{E}}\left(\widehat{\beta}_{j}-\widehat{\beta}^{A;k;\bm{s}}_{j}\right)^{2}}{MSE\left(\widehat{\beta}_{j}\right)}.$ (4.2)
Assuming that the estimation of coefficients is unbiased, the metric defined by Formula (4.2) is motivated by the following equality:
$\mathbb{E}\left[\left(\beta_{j}^{\ast}-\widehat{\beta}_{j}^{\bullet}\right)^{2}\right]$ $=\mathbb{E}\left[\left(\beta_{j}^{\ast}-\widehat{\beta}_{j}+\widehat{\beta}_{j}-\widehat{\beta}_{j}^{\bullet}\right)^{2}\right]=\mathbb{E}\left[\left(\beta_{j}^{\ast}-\widehat{\beta}_{j}\right)^{2}\right]+\mathbb{E}\left[\left(\widehat{\beta}_{j}-\widehat{\beta}_{j}^{\bullet}\right)^{2}\right].$
By construction, $M_{2}\left(A,\bm{s}\right)=1+\delta$, where $\delta>0$. In an ideal situation, $\delta$ is very small, thus meaning that the increase in coefficient value uncertainty is small when synthetic data is used.
We also test all models based on model goodness-of-fit metrics on the unseen dataset $\mathcal{S}_{test}$. Denote by $\mathbf{y}$ and $\widehat{\mathbf{y}}$ the observed and predicted values of the response variable, respectively. Poisson deviance $DEV_{Po}\left(\mathbf{y},\mathbf{\widehat{y}}\right)=2\sum_{i=1}^{n}\left(y_{i}\ln\left(\frac{y_{i}}{\widehat{y}_{i}}\right)-\left(y_{i}-\widehat{y}_{i}\right)\right),$ and $RMSE\left(\mathbf{y},\mathbf{\widehat{y}}\right)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\widehat{y}_{i}\right)^{2}},$ are the model goodness-of-fit metrics, which we use.
Since we are interested in the underlying performance of each approach $A$, all metrics will be averaged across $n_{E}$ different experiments related to $A$. The number of experiments $n_{E}$ should be as large as possible to reduce uncertainty in the evaluation of the algorithm.
Last but not least, we subjectively assess the ease of use of each method. This aspect is often ignored by other studies, as they mainly focus on the performance of the methods.
4.2 Tested methods and formulas
In our case study, we test various methods that can be grouped into four groups: GAN-related methods, MICE-related methods, hybrid methods and other methods. In the description of the methods, the monospaced shorthand used in the results is shown inside the parentheses. Where possible, we use the “off-the-shelf” implementation of a method.
The first group of methods we test are GAN-based, namely:
  • CTGAN (CTGAN) – implementation of Xu et al., (2019) available in the sdv library for Python;
  • CTGAN with AEs (CTGAN_WITH_AE) – a CTGAN enhanced by AEs for encoding-decoding categorical variables;
  • Wasserstein GAN with gradient penalty (COTE_MC_WGAN_GP) – implementation of the method graciously provided by Cote et al., (2020);
  • GAN-based tabulator (JCS_TABULATOR) – tabulator as proposed and implemented by Neves et al., (2022).
The second group of synthetic-data generators we test are the imputation-amputation methods based on MICE. One of the considered approaches is tabulator proposed Neves et al., (2022), but we use MICE as the imputation engine instead of GAIN-based models suggested there. In our “MICE-tabulator”, we have $5$ sequential amputation-imputation iterations, each of each amputates randomly selected $20\%$ of cells and imputes them to generate synthetic data points. The other two methods have a different amputation logic. First, we create a copy of the training data, where we remove $75\%$ of cell values. Second, we append this data with missing values to the original training data and apply MICE to impute the missing values. The data where missing cells were present, is treated as the generated data. In this way, we end up with generated data where 25% of the data comes from the original training dataset and 75% is synthetic. Hereinafter, we refer to this approach as the “MICE method”. The third approach extends the MICE method by the second iteration, in which the cells that were not amputated-imputed in the first iteration are amputated and imputed via MICE. This leads to a data set that is fully synthetic. In summary, the MICE-based methods we consider are
  • MICE method (MICE) – the method described above and implemented using the R-package mice;
  • MICE with all synthetic data (MICE_ALL_SYN) – the modification of the MICE method, where in an additional step the remaining $25\%$ of the original data is imputed via MICE;
  • MICE tabulator (MICE_TABULATOR) – the tabulator setup from Neves et al., (2022) with $5$ iterations and MICE with RFs as the imputation engine.
After testing the above-described groups of methods, we noticed that GAN-based approaches learned the univariate distribution of numeric variables worse than MICE-based methods and had challenges with high-cardinality categorical variables. Due to this finding, we decided to consider in our case studies two hybrid models. One of them combines a CTGAN and a MICE component, namely a CTGAN first generates synthetic data and then $75\%$ of synthetic numeric variable values and imputed amputated and imputed using MICE. Note that also for this CTGAN-MICE model, the data with removed cell values was combined with the original training data to impute the missing numerical column values. The second hybrid model that we tested enhances CTGAN-MICE by using AEs to encode categorical variables before applying CTGAN-MICE and then using clustering with AEs to decode synthetic AE-encoded categorical variables. So, the hybrid approaches we study are:
  • CTGAN with MICE (CTGAN_MICE) – a CTGAN enhanced by MICE for numeric variables;
  • CTGAN with AEs and MICE (CTGAN_WITH_AE_MICE) – a CTGAN enhanced by AEs for categorical variables and MICE for numeric variables.
Lastly, we also test the method proposed in Jamotton and Hainaut, (2024), where variational autoencoders (VAEs) are used for data generation. We use the implementation graciously shared with us by the authors of Jamotton and Hainaut, (2024). The monospaced shorthand for this method is VAE_JAMOTTON.
As described in Subsection 4.1, artificial claim counts are generated for this case study such a way that we know the true relationship between the response variable and the descriptive variables. The following formula for a linear relationship is used:
$\begin{split}f_{lin}(\bm{x})&=-3+0.0075\cdot(\text{VEHICLE\_AGE}-5)+0.0075\cdot(\text{DRIVER\_AGE}-35)\\ &+0.0075\cdot(\text{BONUS\_MALUS}-100)+0.15\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{A,C,E\}}\right)}\\ &-0.5\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{B,F\}}\right)}-0.5\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{4,5,6\}}\right)}\\ &-0.4\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{7,8,9\}}\right)}+0.15\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{12,13,14\}}\right)}\\ &+0.3\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{15\}}\right)}-0.5\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B3,B4,B5\}}\right)}\\ &+0.2\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B10,B11\}}\right)}+0.6\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B13,B14\}}\right)}\\ &+0.45\cdot\mathbbm{1}_{\left(\text{VEHICLE\_GAS}\in\text{\{Diesel\}}\right)}.\end{split}$ (4.3)
Note that the DENSITY and REGION variables available in the freMTPL2freq dataset are not used in generating the claim counts. This is done for the purpose of variable selection testing.
For the interaction formula, two interactions between numeric and categorical variables and two interactions between two categorical variables are added on top of the linear formula:
$f_{int}(\bm{x})$ $=f_{lin}(\bm{x})+0.0015\cdot\text{BONUS\_MALUS}\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{A,B,C\}}\right)}$
$-0.003\cdot\text{BONUS\_MALUS}\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{D,E\}}\right)}$
$+0.015\cdot\text{VEHICLE\_AGE}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{10,11,12,13,14,15\}}\right)}$
$-0.015\cdot\text{VEHICLE\_AGE}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{(4,5\}}\right)}$
$+0.15\cdot\mathbbm{1}_{\left(\text{VEHICLE\_GAS}\in\text{\{Diesel\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{4,5,6,7\}}\right)}$ (4.4)
$-0.25\cdot\mathbbm{1}_{\left(\text{VEHICLE\_GAS}\in\text{\{Regular\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_POWER}\in\text{\{10,11,12,13,14,15\}}\right)}$
$+0.4\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{A,B,D\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B1,B4,B10\}}\right)}$
$+0.2\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{C,D,E\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B2,B6,B11,B12\}}\right)}$
$-0.6\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{A,c,E,F\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B3,B5,B13,B14\}}\right)}$
$-0.3\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{F\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B1,B2,B12\}}\right)}.$
The claim counts simulated using (4.3) and (4.2) lead to the average claim frequency of $5.37\%$ and $6.04\%$ respectively. These numbers are close to the average claim frequencies that we have observed in the industry when modeling MTPL claims.
4.3 Case study results
The case study showcases how different methods for data generation perform across a variety of different performance metrics defined in the previous sections. We do not claim that one particular method is best suited for synthetic data generation, but we assess the strengths and weaknesses of different methods.
With our case study, we evaluated a total of $10$ different approaches to synthetic data generation. However, we display the results only for $9$ of them, as the tested tabulator by Neves et al., (2022) did not generate any number of claims, and, therefore, it could not be fully evaluated.
In total, we have two sets of results, one for the case linear formula (4.3), and one for the case where the interaction formula (4.2) is used. The cases differ in complexity and inter-variable relationships used to generate the artificial number of claims.
For both cases, the methods are assessed based on the model-specific and data-specific metrics. Model-specific metrics are coefficient difference metrics $M_{1}$ (4.1) and $M_{2}$ (4.2), the number of correctly selected variables (using the variable selection method), the number of incorrect or additional variables selected, and lastly Poisson deviance and RMSE of the trained GLM (model with only correct variables). For the interactions case, the number of missing main effects (case where the model has the interaction term but not the corresponding main effect) is also shown.
Data-specific metrics are MAE/MAPE for numeric, categorical and pair-wise combinations of columns and MAE/MAPE for correlation structure within numeric columns of the generated data. Lastly, our subjective rank of the method’s ease of use is also provided.
Note that every synthetic dataset is uniquely generated 5 times ($n_{E}=5$) and all of the following results are presented as simple averages across all generated datasets for a given method (except for $M_{1}$ and $M_{2}$ metrics, as these are already averaged across $n_{E}$). For data-specific metrics, every column is taken as equally important, so the shown metric is an average taken over all columns or column combinations.
The final results for the evaluation of methods are presented in Tables 1 and 2. Each cell shows the rank of the method with respect to the metric shown in the column. Additionally, the table is color matched based on the rank of the method, and in the case of ties, with respect to metric, all cells are given the same highest rank available. Note that in all cases, the ”training” method is shown as the reference value for different metrics.
Generation method M1 M2
Correct
variables
selected
Incorrect
variables
selected
Fit
Poisson
deviance
Fit
RMSE
Pairwise
MAE
Pairwise
MAPE
Correlation
MAE
Correlation
MAPE
Categorical
column
MAE
Categorical
column
MAPE
Numeric
column
MAE
Numeric
column
MAPE
Ease
of use
ranking
training
1
1
(3,15668)
1
1
1
(0,31917)
1
(0,23744)
1
1
1
1
1
1
1
1
1
MICE
2
(33,93479)
2
(33,93813)
1
4
(1,2)
2
(0,32027)
2
(0,2376)
2
(0,00061)
2
(0,33907)
4
(0,00945)
2
(0,27654)
5
(0,00185)
3
(0,02183)
3
(0,0011)
2
(0,27231)
2
CTGAN
3
(41,52358)
3
(41,03633)
1
3
3
(0,32039)
3
(0,23761)
3
(0,00072)
4
(0,44582)
7
(0,01192)
3
(0,33968)
9
(0,00678)
9
(0,11463)
4
(0,00156)
5
(0,37186)
3
MICE_ALL_SYNTHETIC
4
(114,40835)
5
(112,49103)
9
2
(0,8)
4
(0,32396)
4
(0,23812)
4
(0,00107)
5
(0,4497)
5
(0,01031)
4
(0,35079)
6
(0,00232)
4
(0,02677)
2
(0,000995)
4
(0,36295)
3
MICE_TABULATOR
5
(111,53667)
4
(113,48165)
10
(5,6)
8
6
(0,32592)
6
(0,23834)
7
(0,00253)
6
(3,96253)
9
(0,0249)
7
(0,67136)
2
(0,00081)
2
(0,01022)
5
(0,00707)
6
(18,16422)
3
VAE_JAMOTTON
6
(129,11846)
7
(126,94602)
8
(6,6)
4
(1,2)
5
(0,32545)
5
(0,23829)
9
(0,00303)
3
(0,40646)
10
(0,03651)
9
(1,07215)
7
(0,00429)
7
(0,0809)
9
(0,02964)
3
(0,32234)
9
COTE_MC_WGAN_GP
8
(155,80504)
9
(152,74518)
1
8
7
(0,33124)
7
(0,23896)
10
(0,00412)
10
(82,61544)
6
(0,01052)
6
(0,48478)
8
(0,00444)
8
(0,08287)
10
(0,04028)
10
(445,05061)
10
CTGAN_MICE
10
(393,44698)
10
(383,62816)
1
6
(1,4)
9
(0,34255)
9
(0,23977)
8
(0,00292)
8
(8,9032)
2
(0,00543)
5
(0,44177)
9
(0,00678)
9
(0,11463)
7
(0,0147)
8
(42,48769)
6
CTGAN_WITH_AE_MICE
9
(143,03862)
8
(153,73562)
6
(6,8)
6
(1,4)
8
(0,33266)
8
(0,23937)
5
(0,00199)
7
(8,25317)
3
(0,00569)
8
(0,71934)
3
(0,0016)
5
(0,05566)
6
(0,00974)
7
(38,72062)
8
CTGAN_WITH_AE
7
(125,10637)
6
(136,08475)
6
(6,8)
8
10
(0,34588)
10
(0,24254)
6
(0,00244)
9
(17,68936)
8
(0,01736)
10
(1,74036)
3
(0,0016)
5
(0,05566)
8
(0,02108)
9
(83,98836)
6
Table 1: Summary of evaluation results across all tested algorithms when the variable dependence is based on linear formula (4.3).
Looking at Table 1, we can see that in no case does any method outperform training data. The differences between training data and synthetic data can be quite large for example for metric $M_{1}$ or $M_{2}$. However, when it comes to the overall accuracy of the model, the difference in coefficients does not result in a very significant difference in the accuracy of the predictions.
We can see that the best-performing synthetic data generation methods are MICE and CTGAN. They show almost the same accuracy in terms of goodness-of-fit metrics, but MICE is better when it comes to the accuracy of model coefficients. The worst performing methods are CTGAN_MICE, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE. All of these methods not only show worse accuracy in model coefficients but also the model predictions are worse compared to other methods.
Almost all methods show consistency in the correct terms using the variable selection method. However, for all methods the variable selection included additional variables not present in the true model. Here, the worst performing method was MICE_TABULATOR, where automatic variable selection on average failed to include 1.4 correct variables and included, on average, 2 additional incorrect variables (the worst possible result) to the model. The best variable selection results were achieved by models trained on data using the CTGAN method, always selecting the correct variables and including only 1 additional incorrect variable on average.
Similar observations can be seen when it comes to data-specific metrics. Here again, MICE performs the best across all data metrics. From Table 1, we conclude that different methods have different niches. For example, it seems MICE_TABULATOR is much better at categorical column generation, followed closely by CTGAN_WITH_AE and CTGAN_WITH_AE_MICE (since they share categorical variables). This possibly shows that the usage of autoencoders improves the generation of categorical columns. Additionally, we can observe that the addition of MICE to CTGAN-based methods might not be very consistent. MICE helps in terms of numerical column accuracy for CTGAN_WITH_AE_MICE, but makes it worse for CTGAN method. WGAN by Cote et al., (2020) shows overall the worst performance when it comes to data-specific metrics.
As for to the ease of use, MICE is one of the easiest methods to apply, since the implementation in the R package mice is straightforward and requires no additional implementation or data preparation besides the choice of parameters for the function. CTGAN-based methods seem to require more tinkering and also a change of environment to Python since the implementation of the CTGAN interface in R (through the package ctgan) in our experience is neither robust nor intuitive. Addition of autoencoders, although simple in idea, can be quite hard to implement since this requires additional steps with respect to data preparation, setting up and fitting the additional neural networks and finally clustering the generated vectors to get back to the correct categorical labels. The most complex implementations were those provided by Jamotton and Hainaut, (2024) and Cote et al., (2020) since they required the most additional work to make it work properly in our case study.
Generation method M1 M2
Correct
variables
selected
Incorrect
variables
selected
Missing
main effect
variable
Fit
Poisson
deviance
Fit
RMSE
Pairwise
MAE
Pairwise
MAPE
Correlation
MAE
Correlation
MAPE
Categorical
column
MAE
Categorical
column
MAPE
Numeric
column
MAE
Numeric
column
MAPE
Ease
of use
ranking
training
1
1
(0,86047)
1
1
1
1
(0,3394)
1
(0,25286)
1
1
1
1
1
1
1
1
1
MICE
2
(17,6173)
2
(13,62087)
3
(15,4)
6
(1,8)
1
2
(0,34127)
2
(0,25318)
2
(0,00061)
3
(0,34177)
4
(0,00889)
2
(0,30741)
3
(0,002)
3
(0,02169)
3
(0,00139)
3
(0,29525)
2
CTGAN
3
(21,04755)
3
(16,69526)
2
(15,6)
4
(1,6)
6
(0,4)
3
(0,34166)
3
(0,25323)
3
(0,00073)
4
(0,45389)
7
(0,01121)
3
(0,37872)
9
(0,$01$)
9
(0,09023)
4
(0,00175)
5
(0,41926)
3
MICE_ALL_SYNTHETIC
8
(55,7844)
8
(49,90837)
7
(9,4)
7
10
(1,2)
5
(0,34909)
5
(0,25434)
4
(0,00111)
5
(0,4616)
5
(0,00967)
5
(0,38692)
4
(0,00242)
4
(0,02565)
2
(0,00102)
4
(0,3482)
3
MICE_TABULATOR
7
(53,21122)
6
(49,87283)
8
(8,4)
7
6
(0,4)
8
(0,35463)
8
(0,25515)
8
(0,00243)
6
(2,93129)
10
(0,02481)
7
(0,49791)
2
(0,00071)
2
(0,00955)
6
(0,005)
6
(13,22874)
3
VAE_JAMOTTON
6
(55,09532)
7
(49,58059)
10
(7,8)
2
(0,4)
6
(0,4)
7
(0,35376)
7
(0,25501)
6
(0,00203)
2
(0,31078)
9
(0,02237)
8
(0,95891)
6
(0,00523)
7
(0,07658)
7
(0,00784)
2
(0,21804)
9
COTE_MC_WGAN_GP
9
(101,62427)
10
(95,74713)
5
(11,8)
7
3
(0,2)
9
(0,3585)
10
(0,25585)
10
(0,00447)
10
(49,09032)
6
(0,0105)
9
(1,30688)
5
(0,00448)
8
(0,08906)
10
(0,02414)
10
(312,41321)
10
CTGAN_MICE
10
(97,45528)
9
(97,22979)
4
4
(1,6)
3
(0,2)
10
(0,37112)
9
(0,25564)
9
(0,00278)
7
(6,29387)
2
(0,00479)
4
(0,38592)
9
(0,$01$)
9
(0,09023)
9
(0,01193)
7
(30,54492)
6
CTGAN_WITH_AE_MICE
5
(50,98831)
5
(47,65917)
9
7
6
(0,4)
6
(0,35251)
6
(0,25483)
5
(0,00194)
8
(8,69745)
3
(0,00555)
6
(0,47747)
7
(0,00529)
5
(0,06383)
5
(0,00387)
8
(45,25532)
8
CTGAN_WITH_AE
4
(46,39565)
4
(41,8)
6
3
(1,2)
3
(0,2)
4
(0,34617)
4
(0,2539)
7
(0,00207)
9
(19,89066)
8
(0,02117)
10
(1,79309)
7
(0,00529)
5
(0,06383)
8
(0,00852)
9
(103,25122)
6
Table 2: Summary of evaluation results across all tested algorithms when the variable dependence is based on interaction formula (4.2).
Table 2 shows very similar results to what we saw from in Table 1. No method is able to beat the results we observe for training data. However, here we can see that the model accuracy metrics, even for the best methods, are further away than in the linear case. In other words, this indicates that the generator fidelity depends on the underlying data dependence structure.
From model-specific results, we see that coefficient distance metrics $M_{1}$ and $M_{2}$ have much smaller values compared to the linear case. The best methods based on coefficient difference metrics are again MICE and CTGAN, and the worst method seems to be COTE_MC_WGAN_GP and CTGAN_MICE. Notably, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE show a big uplift in rankings while additional MICE-based methods lose ranks compared to the linear case. This possibly shows that MICE is less suited for data structures with inherent dependence, but overall performance still seems to be the best across the tested approaches.
Model goodness-of-fit metrics reflect very similar patterns as model coefficient metrics. However, MICE_ALL_SYNTHETIC shows a higher rank with respect to the goodness-of-fit metrics. The Poission deviance for it is notably smaller (0.00554) than that for MICE_TABULATOR.
The variable selection shows, however, much more varied results. Again, MICE and CTGAN methods show best results, where variable selection, on average, selected 15.4 and 15.6 correct terms. Beyond these two methods, the variable selection in the next best method selected, on average, 12 correct terms. Variable selection for VAE_JAMOTTON got the worst result, on average, selecting only 7.8 correct terms to the model. We can also see that for almost all methods, variable selection added additional incorrect terms, where only VAE_JAMOTTON picked on average 0.4 additional incorrect terms. These combined facts indicate that using VAE_JAMOTTON results in a small/lean model, which is good for the linear case, but worse for the interaction case.
From the data-specific results, we can observe almost the same rankings as we did for the linear case. This is expected since we did not introduce any additional inter-variable relationships besides the interaction relationship between response and the two covariates. The biggest difference is the fact that autoencoders did not improve the MAE/MAPE of categorical columns as significantly for CTGAN method as they had done in the linear case study. MICE is again clearly the leader when it comes to data-specific metrics, achieving top ranks in all metrics. However, analyzing synthetic data on a more granular level, we noticed that MICE-based methods can generate numeric values that ignore business constraints, e.g., generating observations with a driver age below $18$. Imposing such constraints may be investigated in a follow-up paper.
Figure 2: $M_{1}$ metric results as a function of synthetic data proportion. ”All synthetic” represents the cases where only synthetic data is used and are treated as ”asymptotic” estimation of the shown relationship. The ”All synthetic” metric values are scaled 1:2 (right side axis) compared to the other presented values (left side axis). The dotted lines represent the corresponding asymptotic trends.
Having discussed data generation fidelity, we now focus on data augmentation. For this, we look at $M_{1}$ metric depending on the proportion of synthetic data added. We observed that as we increase the proportion of additional synthetic data compared to the original training data, the value for $M_{1}$ metric increases. The only case where we could observe a decrease in a metric was for coefficient distance metrics $M_{1}$ for the method MICE with an additional synthetic data proportion of $0.2$. In this case, the augmented data achieved the $M_{1}$ metric value of $0.78717$ while the corresponding training data metric was $0.86047$. Since this is the only result where this happened, it is not enough to say that augmentation of the original freMTPL2freqdata might help when it comes to better accuracy of coefficients.
The general relationship between the increase of additional synthetic data proportion and metric value can be seen in Figure 2. The shown figure is for metric $M_{1}$. As the synthetic data proportion increases, we can see a clear increasing trend for $M_{1}$ metric value for the given method. Based on the way $M_{1}$ metric is calculated and how GLM relies on the underlying data to compute the coefficient values, we estimate that if we keep on increasing the proportion of additional synthetic data, then we will asymptotically reach the metric values which we can compute for case where we use only synthetic data. These pure synthetic data results are shown on the right side of the plot.
5 Conclusion
Our paper demonstrates that the amputation-imputation framework based on MICE can be effectively used within an actuarial context. This framework is competitive for generating synthetic ratemaking data when compared to more sophisticated deep generative models such as CTGANs or VAEs.
In our case studies based on the open-source freMTPL2freq dataset, synthetic data generated using MICE with random forests as the imputation engine shows, on average, the best properties in terms of modeling efficacy and achieves high ranks in model accuracy and coefficient difference metrics. In general, several of the $9$ tested methods for synthetic data generation seem to have their niches for specific use cases.
Having subjectively evaluated the ease of use for each tested method, we identify MICE as the most user-friendly approach, since it requires the least amount of additional work to get started, which is an important practical aspect for actuaries. Custom deep generative models like WGAN by Cote et al., (2020) or VAE by Jamotton and Hainaut, (2024) seem to require significantly more expertise and tuning to get good results.
Our case study also shows that generic data augmentation does not generally improve the performance of GLMs trained on the real data augmented with the synthetic one. We observe that an increase in the proportion of synthetic data compared to the original training data leads to an approximately linear increase in the metrics that measure the difference between coefficients of two GLMs, e.g., the true GLM vs a GLM estimated on augmented data. We have noticed only one case where data augmentation improved the estimation of the true GLM coefficients.
We leave several aspects for future research. For example, it may be interesting to investigate how to integrate business constraints into the amputation-imputation framework based on MICE, e.g., linear constraints on numerical variables or hierarchy among categorical variables. Another relevant topic for further research may be the privacy of synthetic datasets generated with the help of MICE.
Acknowledgments
Yevhen Havrylenko acknowledges the financial support by the PRIME program (link) of the German Academic Exchange Service (DAAD, https://ror.org/039djdh30) with funds from the German Federal Ministry of Education and Research (BMBF). Yevhen Havrylenko acknowledges the support of the University of Copenhagen and Johannes Kepler University Linz, where he conducted parts of this research project. Meelis Käärik and Artur Tuttar acknowledge the support of the Estonian Research Council grant PRG1197.
The authors of this publication express their gratitude to Raul Kangro and Märt Möls for their insights, ideas, and discussion related to the article. The authors are grateful to the participants of the 28th International Congress on Insurance: Mathematics and Economics.
Declaration of generative AI and AI-assisted technologies
The authors used ChatGPT (free version) and Gemini (Flash and Pro) models to improve the language of the manuscript. After using those large language models, the authors reviewed and edited the content as needed and take full responsibility for the content of the article.
Declaration of competing interest
The authors declare no conflict of interest.
Appendix A Appendix A
In this appendix, we explain how deterministic AEs are combined with synthetic data generators (e.g., CTGAN) to potentially improve the quality of synthetic categorical variables.
Before training a generation method $GM_{A}$, we perform the following steps for each categorical variable $x_{j}$:
  1. 1.
    represent $x_{j}$ as $\bm{x}_{j}^{cat}$ via one-hot encoding;
  2. 2.
    train a separate AE (see Section 3.3), which finds a low-dimensional representation (encoding) $\varphi_{j}(\bm{x}_{j}^{cat})$ of $\bm{x}_{j}^{cat}$;
  3. 3.
    in $\mathcal{S}_{train}$, replace $x_{j}$ by $\varphi_{j}(\bm{x}_{j}^{cat})$.
Since the generation method $GM_{A}$ generates data in the same form as it is trained on, the synthetic data will have a low-dimensional numeric representation of the categorical variables. This is an issue since the meaning of the categorical variables is not apparent. To resolve this, we perform the following steps for each vector of features corresponding to a categorical variable (in its quasi-encoded representation) in the synthetic dataset:
  1. 1.
    use Nearest Centroid Classification to identify for this vector its the closest class; the centroid of each class is determined by the encoding (given by the respective AE fitted before training $GM_{A}$) of the respective category of the original categorical variable;
  2. 2.
    replace this vector by the one-hot encoded representation of the category that corresponds to the centroid of the class to which the generated vector belongs; this one-hot encoded representation can be recovered from the centroid using the decoding part of AE.

### v2

@@
-Amputation-imputation based generation of synthetic tabular data for ratemaking
+Synthetic data for ratemaking: imputation-based methods vs adversarial networks and autoencoders
 Abstract
@@
 In this paper, we explore synthetic-data generation as a potential solution to these issues.
-In addition to discussing generative methods previously studied in the actuarial literature, we introduce to the insurance community another approach based on Multiple Imputation by Chained Equations (MICE).
-We present a comparative study using an open-source dataset and evaluating MICE-based models against other generative models like Variational Autoencoders and Conditional Tabular Generative Adversarial Networks.
+In addition to generative methods previously studied in the actuarial literature, we explore and benchmark another class of approaches based on Multivariate Imputation by Chained Equations (MICE).
+In a comparative study using an open-source dataset, MICE-based models are evaluated against other generative models like Variational Autoencoders and Conditional Tabular Generative Adversarial Networks.
 We assess how well synthetic data preserves the original marginal distributions of variables as well as the multivariate relationships among covariates.
-We also investigate the consistency between Generalized Linear Models (GLMs) trained on synthetic data with GLMs trained on the original data.
-Furthermore, we assess the ease of use of each generative approach and study the impact of augmenting original data with synthetic data on the performance of GLMs for predicting claim counts.
-Our results highlight the potential of MICE-based methods in creating high-quality tabular data while being more user-friendly than the other methods.
-Keywords: synthetic actuarial data, data augmentation, multiple imputation by chained equations, generative adversarial networks, autoencoders
+The consistency between Generalized Linear Models (GLMs) trained on synthetic data with GLMs trained on the original data is also investigated.
+Furthermore, we assess the ease of use of each generative approach and study the impact of generically augmenting original data with synthetic data on the performance of GLMs for predicting claim counts.
+Our results highlight the potential of MICE-based methods in creating high-fidelity tabular data while offering lower implementation complexity compared to deep generative models.
+Keywords: synthetic actuarial data, data augmentation, multivariate imputation by chained equations, generative adversarial networks, autoencoders
 1 Introduction
@@
 Even if insurers and reinsurers may have enough data for some products or business lines, the companies rarely publicly share their data due to concerns related to data privacy, security, and their competitive advantage.
-As a result, the number of open-access actuarial datasets is quite small (see, e.g., Gabrielli and Wüthrich, (2018), Cote et al., (2020), Segura-Gisbert et al., (2024)).
+As a result, the number of open-access actuarial datasets is quite small (see, e.g., Gabrielli and Wüthrich (2018), Côté et al.
+(2025), Segura-Gisbert et al.
+(2024)).
 Furthermore, the companies themselves may also face a lack of data, e.g., when they enter a new market or release a new (re-)insurance product.
-To increase the amount of actuarial data, one can generate synthetic data, which can be used as a standalone dataset or as an enlargement to the existing dataset.
+To increase the amount of actuarial data, one can generate synthetic data, which can be used as a standalone dataset or as an enlargement to an existing dataset.
 In the former case, synthetic data can be publicly shared and used for research purposes (e.g., testing new models and algorithms), if it has the statistical properties of the real data and does not expose sensitive information.
@@
 We will focus on the latter type, as one of our aims is to develop an actuarial data generator that can be easily used “out-of-the-box”.
-Furthermore, the insurance application we target is general ratemaking (without telematics), where, usually, datasets are tabular with both numeric and categorical variables and temporal independence of observations is assumed.
+Furthermore, the insurance application we target is general ratemaking, where, usually, datasets are tabular with both numeric and categorical variables and temporal independence of observations is assumed.
 The generation of synthetic data without the need for distributional assumptions gained much popularity due to the advancement of deep generative models such as generative adversarial networks (GANs) and variational autoencoders (VAEs).
-To our knowledge, only three papers investigated the use of such models in the context of ratemaking: Kuo, (2020), Cote et al., (2020), Jamotton and Hainaut, (2024).
+To our knowledge, only three papers investigated the use of such models in the context of ratemaking: Kuo (2020), Côté et al.
+(2025), Jamotton and Hainaut (2024).
 We discuss these articles in more detail in the literature overview below and mention for now only aspects that strongly motivate us to conduct further research on this topic.
@@
 Second, the benefit of augmenting original data with synthetic data is not addressed in those papers.
-Our contributions.
-First, we introduce to actuarial science several amputation-imputation-based approaches to data generation.
-These techniques are based on Multiple Imputation by Chained-Equations (MICE) with Random Forests (RFs) as imputation models.
-According to our tests, these approaches can be more suitable for “out-of-the-box” usage in practice than the generative models studied in Kuo, (2020), Cote et al., (2020), Jamotton and Hainaut, (2024), while having a comparable performance.
-Second, using a publicly available French Motor Third-Part Liability dataset called freMTPL2freq (Dutang and Charpentier,, 2019) as a training dataset, we compare $9$ approaches to ratemaking data generation, including the existing approaches (Kuo,, 2020; Cote et al.,, 2020; Jamotton and Hainaut,, 2024), their modifications that address the challenge of generative adversarial networks (GANs) in dealing with high-cardinality categorical variables, and the newly introduced amputation-imputation-based methods.
-We analyze the quality of synthetic datasets by studying two aspects: the similarity between the distribution of the original data and of synthetic data, the ability of methods to capture structural dependence among covariates and the response variable.
-Furthermore, we assess the impact of augmenting original data with synthetic data on the performance of GLMs for predicting claim frequency.
-Our main results are threefold.
-First, we illustrate that the MICE-RF approach is a competitive method for generation of synthetic ratemaking data in comparison to deep generative models such as a VAE and a CTGAN.
-The important practical advantage of the MICE-RF algorithm is its ease of use, given the existing packages in $\mathrm{R}$ and the versatility of RFs in dealing with variables of different types without much pre-processing.
-Second, we show that modifying a CTGAN with an AE for categorical-variable pre-processing improves the quality of synthetic high-cardinality categorical variables but deteriorates the performance of this hybrid approach in terms of other criteria.
-In general, a MICE-RF approach and a CTGAN without AE seem to be the most balanced in terms of the performance metrics we consider.
-Third, our case studies indicate that for our experimental data sets based on freMTPL2freq data augmentation does not improve the performance of GLMs trained on the original data merged with synthetic one.
 Literature overview.
-Kuo, (2020) pioneered the use of conditional tabular GANs for generating actuarial datasets, focusing on ratemaking and lapse modeling.
-In his experiments with freMTPL2freq dataset, the researcher showed that generalized linear models (GLMs) trained on synthetic data achieved performance comparable to GLMs trained on real data, and the distributions of both continuous and categorical variables in the synthetic datasets closely mirrored the real datasets.
-However, the relativities (how much each level of a categorical variable affects the predicted outcome compared to a reference level) of GLM models constructed using synthetic datasets do not show the same patterns as those of models created using real datasets.
+Kuo (2020) pioneered the use of conditional tabular GANs for generating actuarial datasets, focusing on ratemaking and lapse modeling.
+In his experiments with the freMTPL2freq dataset, the researcher showed that generalized linear models (GLMs) trained on synthetic data achieved performance comparable to GLMs trained on real data, and the distributions of both continuous and categorical variables in the synthetic datasets closely mirrored the real datasets.
+However, the relative effects (how much each level of a categorical variable affects the predicted outcome relative to a reference level) of GLM models constructed using synthetic datasets do not show the same patterns as those of models constructed using real datasets.
 Because of the observed instability in the parameters of a GLM trained on synthetic data, developing a rating plan based on it might not be the best approach.
-Inspired by Kuo, (2020), Cote et al., (2020) investigated three GAN-based models for generating synthetic insurance data: Multi-Categorical Wasserstein GAN with Gradient Penalty (MC-WGAN-GP), Conditional Tabular GAN (CTGAN), and Mixed Numerical and Categorical Differentially Private GAN (MNCDP-GAN).
+Inspired by Kuo (2020), Côté et al.
+(2025) investigate three GAN-based models for generating synthetic insurance data: Multi-Categorical Wasserstein GAN with Gradient Penalty (MC-WGAN-GP), Conditional Tabular GAN (CTGAN), and Mixed Numerical and Categorical Differentially Private GAN (MNCDP-GAN).
 In a case study using the freMTPLfreq data, they find that MC-WGAN-GP performs best in replicating univariate distributions and the observed claim frequency for each category of categorical variables, followed by CTGAN, and then MNCDP-GAN.
@@
 The authors conclude that neither of the tested models performs well across all key performance indicators (KPIs) with a high level of training data confidentiality.
-As an alternative to GAN-based models, Jamotton and Hainaut, (2024) propose a Variational Autoencoder (VAE) model for the generation of insurance data.
+As an alternative to GAN-based models, Jamotton and Hainaut (2024) propose a Variational Autoencoder (VAE) model for the generation of insurance data.
 Their approach features a quantile transformation for continuous variables to address multimodality.
 In their VAE, the researchers use a reconstruction loss function that consists of two parts: cross-entropy for categorical variables and MSE for continuous variables.
-In a case study using a data set from the homepage of the book Ohlsson and Johansson, (2010), the researchers find that the marginal distributions of the real datasets and those of the synthetic dataset are almost the same, as supported by visual inspection, Kolmogorov-Smirnov tests for continuous variables, and Chi-Squared tests for categorical variables.
+In a case study using a dataset from the homepage of the book Ohlsson and Johansson (2010), the researchers find that the marginal distributions of the real datasets and those of the synthetic dataset are almost the same, as supported by visual inspection, Kolmogorov-Smirnov tests for continuous variables, and Chi-Squared tests for categorical variables.
 Based on Pearson and Spearman correlations for continuous variables and contingency scores based on contingency tables for categorical variables, the authors conclude that synthetic data preserves pairwise relationships among variables reasonably well.
 Training a GLM on synthetic data generated by their VAE and noticing that the respective Poisson deviance is almost the same as the one for a GLM trained on the original data, the scientists conclude that synthetic data can substitute real data in actuarial analyses.
-Fundamentally different data generation methods are those based on data imputation and have not yet been explored in the actuarial science context.
-The main idea behind those methods is to sequentially “amputate” parts of data and then impute it with a suitable statistical learning model.
-To our knowledge, this idea was first introduced by D.B.
-Rubin, who first developed a data imputation approach called MICE in the 1970s (cf.
-Rubin, (1977)) then proposed to use this approach also to create (fully or partially) synthetic datasets in Rubin, (1993).
-We refer readers interested in the subsequent development of MICE-based generation of synthetic data to Drechsler, (2011) and emphasize two aspects.
-First, this approach gained popularity in various scientific domains (e.g., see Volker and Vink, (2021) for medicine), yet it seems to have been overseen by the actuarial community.
-Second, the imputation in this framework is done on a variable-by-variable (column-by-column) basis, and the imputation model can be flexibly chosen from linear regression to sophisticated models.
-A related idea of generating artificial data using imputation appeared in Neves et al., (2022), where the authors propose a so-called tabulator framework.
-In their approach, they propose to randomly (not on variable-by-variable basis) select data entries to be amputated and then imputed by a so-called back-end engine for imputation.
-As a model driving this imputation engine the authors propose to use one of the three methods that can be seen as modifications of Generative Adversarial Imputation Networks (GAINs, originally proposed by Yoon et al., (2018)): Slim GAIN (SGAIN), Wasserstein Slim GAIN with Clipping Penalty (WSGAIN-CP), Wasserstein Slim GAIN with Gradient Penalty (WSGAIN-GP).
-According to the researchers, tabulator has a comparable performance (tested on 10 datasets from various fields like healthcare, economics, and computer science) with CTGAN that, in turn, is better than medGAN, VEEGAN, and table-GAN.
+Fundamentally different data generation methods are those based on multivariate imputation of data.
+While less explored in actuarial science, these methods are well-established in statistical disclosure control.
+The conceptual foundation was laid by Rubin (1993), who proposed treating the creation of synthetic data as a missing data problem: synthetic values are drawn from the posterior predictive distribution of the original data.
+To obtain the predictive distribution of the original data, it is common to use one of the two approaches: Sequential Regression (SR) or Multivariate Imputation by Chained Equations (MICE).
+The SR approach approximates the joint multivariate distribution of data by explicitly factorizing it into a strictly ordered chain of univariate conditional densities (i.e., 1st variable, then 2nd variable given the 1st variable, and so on).
+Reiter (2005) demonstrates that classification and regression trees (CART) can be effectively integrated into the SR framework.
+Building on that paper, Caiola and Reiter (2010) show that using Random Forests (RFs) as imputation models has multiple benefits for creating synthetic datasets using SR: the synthesizer requires minimal fine-tuning, naturally handles variables of any type, scales well to high-dimensional data, and automatically captures non-linear relationships.
+This synthetic data generation methodology is implemented in the R-package synthpop, see Nowok et al.
+(2016).
+SR approach reduces the synthesis task to a series of unidirectional regression steps according to a user-defined “synthesis order” (visit.sequence parameter in synthpop package), e.g., first Age, then Income given Age, then Family Status given Income and Age, and so on).
+As mentioned in Caiola and Reiter (2010) and Drechsler and Reiter (2011), there is no mathematical theory underpinning the ordering of variables for generation, so different choices of the sequence could produce synthetic datasets of different quality.
+Once the user fixes the synthesis sequence, earlier variables can influence the generation of the later ones, but not the other way around, which is why the dependence structure among synthetically generated variables may not be close enough to the true dependence structure and may lead to biased estimation of multivariate relationships (Reiter, 2005).
+Since ratemaking data can contain variables with endogenous, bidirectional dependencies (e.g., the interplay between coverage limits, deductibles, and vehicle value) that lack a natural synthesis order, we do not explore this method in our paper and focus on MICE, which is independent of the order in which variables are generated.
+Developed at the end of the nineties, MICE operates by modeling the conditional density of each variable given all other variables, which is why it is also called the Fully Conditional Specification approach.
+This is an important difference to the SR methodology in which every single variable is imputed/generated conditional on a model trained on only the variables located before the current variable in the visiting sequence.
+Similarly to SR, MICE handles diverse variable types without explicitly defining a high-dimensional joint distribution, c.f., Drechsler (2011), but it offers more flexibility and is easier to use in practice because it does not require any ordering of variables for data generation.
+Instead, MICE utilizes an iterative Gibbs sampling scheme.
+By cycling through the covariates multiple times, the algorithm relies on the asymptotic convergence of the Markov Chain to reach a stationary distribution that reflects the complex, multidimensional dependence structure of the data.
+This approach to data imputation is implemented in the R-package mice, see van Buuren and Groothuis-Oudshoorn (2011).
+As with SR, MICE can also be used for synthetic data generation, which is demonstrated by Volker and Vink (2021) for a medicine-related dataset with $9$ variables and about $750$ observations.
+The researchers highlighted that MICE with CART as the imputation model produces indistinguishable synthetic records that support unbiased, valid statistical inferences, and offer a user-friendly solution for securely sharing and analyzing sensitive data.
+Although SR and MICE methodologies for synthetic data generation have been gaining popularity and credibility in the medical and social sciences, they seem to have been largely overlooked by the actuarial community in favor of deep generative models described at the beginning of our literature overview.
+A related idea of generating artificial data using imputation appeared in Neves et al.
+(2022), where the authors propose a so-called tabulator framework.
+In their approach, they propose to randomly (not on variable-by-variable basis) select data entries to be set to missing and then be imputed by a so-called back-end engine for imputation.
+As a model driving this imputation engine, the authors propose to use one of the three methods that can be seen as modifications of Generative Adversarial Imputation Networks (GAINs, originally proposed by Yoon et al.
+(2018)): Slim GAIN (SGAIN), Wasserstein Slim GAIN with Clipping Penalty (WSGAIN-CP), Wasserstein Slim GAIN with Gradient Penalty (WSGAIN-GP).
+According to the researchers, tabulator has a comparable performance (tested on 10 datasets from various fields like healthcare, economics, and computer science) with CTGAN that, in turn, is better than medical Generative Adversarial Network (medGAN), Variational Encoder Enhanced Generative Adversarial Network (VEEGAN), and Table Generative Adversarial Network (table-GAN).
 The researchers conclude that the tabulator preserves the dependence structure of the original dataset better than CTGAN, but both approaches struggle when generating synthetic data from original imbalanced data.
 It is also stated that synthetic data from tabulator improves the performance of regression/classification tasks only marginally.
-Readers interested in the literature on data generation beyond traditional ratemaking are referred to So et al., (2021) and So and Jeong, (2022) for telematics data, to Gabrielli and Wüthrich, (2018), Avanzi et al., (2021), Avanzi et al., (2023), Bauwelinckx et al., (2024) for reserving data, to Campo and Antonio, (2023) for auto-insurance fraud-detection data, and to Chapter 10 in Wuthrich et al., (2025) for a brief introduction to generative modeling tools for actuaries.
-We refer readers to Fonseca and Bacao, (2023) and Brophy et al., (2023) for a systematic review of the literature on the GAN-based generation of tabular data and on the GAN-based generation of time series data, respectively.
-For an overview comparing deep generative models, including but not limited to VAEs and GANs, we direct readers to Bond-Taylor et al., (2022).
+Recent comprehensive reviews of literature on synthetic data generation indicate that deep generative models do not consistently outperform established methods for tabular data.
+Reiter (2023) observes that statistical agencies, such as the U.S.
+Census Bureau, continue to rely on sequential conditional modeling for major data products, citing CART as a “primary engine” or “leading contender” for synthesis tasks.
+Furthermore, Drechsler and Haensch (2024) report that despite the great success of GANs in computer science, comparative evaluations reveal that sequential regression based on CART frequently offers the highest utility but also the highest disclosure risk.
+Readers interested in the literature on insurance data generation beyond traditional ratemaking are referred to So et al.
+(2021) and So and Jeong (2022) for telematics data, to Gabrielli and Wüthrich (2018), Avanzi et al.
+(2021), Avanzi et al.
+(2023), Bauwelinckx et al.
+(2024) for reserving data, to Campo and Antonio (2023) for auto-insurance fraud-detection data, and to Chapter 10 in Wuthrich et al.
+(2025) for a brief introduction to generative modeling tools for actuaries.
+We refer readers to Fonseca and Bacao (2023) and Brophy et al.
+(2023) for a systematic review of the literature on the GAN-based generation of tabular data and on the GAN-based generation of time series data, respectively.
+For an overview comparing deep generative models, including but not limited to VAEs and GANs, we direct readers to Bond-Taylor et al.
+(2022).
+Our contributions.
+First, we benchmark the utility and practicality of several amputation-imputation-based approaches to data generation for actuarial ratemaking.
+These techniques are based on Multivariate Imputation by Chained-Equations (MICE) with Random Forests (RFs) as imputation models.
+According to our tests, these approaches can be more suitable for “out-of-the-box” usage in practice than the generative models studied in Kuo (2020), Côté et al.
+(2025), Jamotton and Hainaut (2024), while having a comparable performance.
+Second, using a publicly available French Motor Third-Part Liability dataset called freMTPL2freq (Dutang and Charpentier, 2019) as a training dataset, we compare $10$ approaches to ratemaking data generation, including the existing approaches (Kuo, 2020; Côté et al., 2025; Jamotton and Hainaut, 2024), their modifications that address the challenge of GANs in dealing with high-cardinality categorical variables, and the newly introduced amputation-imputation-based methods.
+We analyze the quality of synthetic datasets by studying two aspects: the similarity between the distribution of the original data and of synthetic data, the ability of methods to capture structural dependence among covariates and the response variable.
+Furthermore, we assess the impact of generically augmenting original data with synthetic data on the performance of GLMs for predicting claim frequency.
+Our main results are threefold.
+First, we illustrate that the MICE-RF approach is a competitive method for generation of synthetic ratemaking data in comparison to deep generative models such as a VAE and a CTGAN.
+The important practical advantage of the MICE-RF algorithm is its ease of use, given the existing packages in R and the versatility of RFs in dealing with variables of different types without much pre-processing.
+Second, we show that modifying a CTGAN with an AE for categorical-variable pre-processing improves the quality of synthetic high-cardinality categorical variables but deteriorates the performance of this hybrid approach in terms of other criteria.
+In general, MICE-RF-based approaches are the best-performing in terms of most performance metrics we consider in our experiments.
+Third, our case studies indicate that for our experimental datasets based on the freMTPL2freq, data augmentation does not improve the performance of GLMs trained on the original data merged with synthetic one.
 Structure of the paper.
@@
 Third, we explain what is meant by generating synthetic data and state two research questions that are answered in the subsequent sections.
-We denote a tabular dataset by $\mathcal{S}:=\left\{\bm{s}_{i}\right\}_{i=1,\dots,n_{ROD}}$, where $n_{ROD}$ is the number of rows in the original data (number of observations), $\bm{s}_{i}=(\bm{x}_{i},v_{i},y_{i})$ is a tuple with three elements such that $\bm{x}_{i}=(x_{i;1},\dots,x_{i;j},\dots,x_{i;n_{C}})^{\top}$ is a $n_{C}$-dimensional vector of observed values of covariates (independent variables), $n_{C}\in\mathbb{N}$, is the transpose operator, $v_{i}\in(0,1]$ is an exposure-to-risk (proportion of a year during which the observation $i$ was tracked) and $y_{i}$ is an observed value of an one-dimensional response variable (dependent variable) for anobservation $i=1,\dots,n_{ROD}$.
+We denote a tabular dataset by $\mathcal{S}:=\left\{\bm{s}_{i}\right\}_{i=1,\dots,n_{ROD}}$, where $n_{ROD}$ is the number of rows in the original data (number of observations), $\bm{s}_{i}=(\bm{x}_{i},v_{i},y_{i})$ is a tuple with three elements such that $\bm{x}_{i}=(x_{i;1},\dots,x_{i;j},\dots,x_{i;n_{C}})^{\top}$ is a $n_{C}$-dimensional vector of observed values of covariates (independent variables), $n_{C}\in\mathbb{N}$, is the transpose operator, $v_{i}\in(0,1]$ is an exposure-to-risk (proportion of a year during which the observation $i$ was tracked) and $y_{i}$ is an observed value of a one-dimensional response variable (dependent variable) for an observation $i=1,\dots,n_{ROD}$.
 A tuple $\bm{s}_{i}$ can be interpreted as an insurance policy snippet, $\bm{x}_{i}$ – as a vector of policy characteristics, $v_{i}$ – as the fraction of a year during which a policy snippet $i$ was active, and $y_{i}$ – as the observed number of claims (or the monetary loss associated with a claim in the case of severity modeling).
@@
 The vector $\bm{x}$ can have both categorical and numerical components.
-We denote by $n_{CC}$ the number of categorical covariates in $\bm{x}$ and by $n_{NC}=n_{C}-n_{CC}$ the number of numerical covariates in $\bm{x}$.
+We denote the number of categorical covariates in $\bm{x}$ by $n_{CC}$ and the number of numerical covariates in $\bm{x}$ by $n_{NC}$: $n_{NC}=n_{C}-n_{CC}$.
 Before fitting/training any model on the raw data, it is usually pre-processed to a suitable format, depending on the model to be used later.
@@
 In ratemaking, it is common to model the link between $y$ and $(\bm{x},v)$ via a generalized linear model (GLM).
-GLMs were introduced in Nelder and Wedderburn, (1972) and became very popular in insurance, where they are routinely used in various business lines such as claim management, underwriting, etc.
+GLMs were introduced in Nelder and Wedderburn (1972) and became very popular in insurance, where they are routinely used in various business lines such as claim management, underwriting, etc.
 In our paper, we will only discuss the basics of a Poisson GLM, which is frequently used to predict insurance claim counts and which we use in numerical studies.
-For a deep dive into the theory and practice of GLMs in insurance readers are referred to Chapter 4 of Denuit and Trufin, (2019) and Chapter 5 of Wüthrich and Merz, (2023).
+For a deep dive into the theory and practice of GLMs in insurance readers are referred to Chapter 4 of Denuit and Trufin (2019) and Chapter 5 of Wüthrich and Merz (2023).
 Let $Y$ be a random variable representing a response variable (e.g., number of insurance claims per year) and $y$ be its realization (observation).
@@
 To consider an interaction with an entire categorical variable, a sum of different (2.3) (one for each category in the respective categorical variable, except for the dummy one) must be added to the right-hand side of (2.2).
-For more information on modeling interactions in GLMs, see, e.g., Fox and Weisberg, (2018), or Section 4.2.5 in Denuit and Trufin, (2019).
-Let us denote by $\bm{\beta}^{\ast}$ the true value of $\bm{\beta}$.
+For more information on modeling interactions in GLMs, see, e.g., Fox and Weisberg (2018), or Section 4.2.5 in Denuit and Trufin (2019).
+Let us denote the true value of $\bm{\beta}$ by $\bm{\beta}^{\ast}$.
 In practice, $\bm{\beta}^{\ast}$ is not known and must be estimated from collected data.
@@
 To address the aspect of data quantity and quality, an actuary can generate a synthetic dataset $\mathcal{S}^{A}$ of a desirable size and reasonable quality.
-In this situation, two natural questions arise: one is related to the usage of synthetic data as a standalone set, and the other one is related to working with the original dataset that is augmented with a part of the synthetic dataset.
-First, how well does $\mathcal{S}^{A}$ preserve (from training data $\mathcal{S}_{train}$) the marginal distributions of variables as well as the multivariate relationships among covariates and how consistent is a GLM trained on $\mathcal{S}^{A}$ with a GLM fitted on $\mathcal{S}_{train}$?
+In this situation, two natural questions arise: one is related to the usage of synthetic data as a standalone set, and the other is related to working with the original dataset that is augmented with a part of the synthetic dataset.
+First, how well does $\mathcal{S}^{A}$ preserve (from training data $\mathcal{S}_{train}$) the marginal distributions of variables as well as the multivariate relationships among covariates, and how consistent is a GLM trained on $\mathcal{S}^{A}$ with a GLM fitted on $\mathcal{S}_{train}$?
 Second, which advantages and disadvantages does training GLMs on an augmented dataset $\mathcal{S}_{train}\cup\mathcal{S}^{A}$ have in terms of GLM’s predictive performance and the estimates of $\bm{\beta}^{\ast}$ (e.g., $\widehat{\bm{\beta}}(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ and/or $\widehat{\widehat{\bm{\beta}}\,}\!(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ are closer to the true $\bm{\beta}^{\ast}$)?
@@
 Third, we summarize deterministic and variational autoencoders (AEs).
-3.1 Multiple imputation by chained equations
-According to van Buuren, (2018), multiple imputation by chained equations (MICE) is one of the best approaches to dealing with missing data.
+3.1 Multivariate imputation by chained equations
+According to van Buuren (2018), multivariate imputation by chained equations (MICE) is one of the best approaches to dealing with missing data.
 MICE is also known as fully conditional specification, sequential regression multivariate imputation, or simply chained equations.
@@
 This process of iterative drawing can be seen as a Gibbs sampler that will converge to the draws from the theoretical joint distribution of data, should this joint distribution exist (cf.
-Raghunathan et al., (2001)).
+Raghunathan et al.
+(2001)).
 In practice, the assumption about the existence of the joint distribution is often implicitly made.
-Borrowing notation from Drechsler, (2011), we summarize the MICE algorithm:
+Borrowing notation from Drechsler (2011), we summarize the MICE algorithm:
   1. 1.
@@
   6. 6.
-    Iterate Steps 2-4 for several cycles to update imputations until the convergence or the maximal number $n_{MI}$ of MICE iterations is reached.
+    Iterate Steps 2-5 for several cycles to update imputations until the convergence or the maximal number $n_{MI}$ of MICE iterations is reached.
   7. 7.
@@
 Figure 1 visually illustrates the original MICE approach.
-[…]
+Figure 1: Schematic representation of MICE.
+$\bm{Y}^{(t-1)}_{-j}:=\left(Y^{(t)}_{1},\dots,Y^{(t)}_{j-1},Y^{(t-1)}_{j+1},\dots,Y^{(t-1)}_{p}\right)$, $Y^{(t)}_{j,mis}$ denotes the elements of the column corresponding to $Y^{(t)}_{j}$ that must be imputed ($\Bigl$(i.e., missing in the column corresponding to $Y^{(0)}_{j}\Bigr)$, $Y^{(t)}_{j,obs}$ denotes the elements in the respective column that were originally observed in $Y^{(0)}_{j}$, $f_{j}$ is an imputation model (e.g., a GLM, a random forest, etc.), $\bm{\theta}^{(t)}_{j}$ is the vector of parameters of $f_{j}$ in MICE iteration $t$.
 As for Step 7, the examples of factors contributing to the variability of the imputed datasets are: initial imputation of the missing data in Step 1 (mean, median, random draw from observed values, etc.), specification of the imputation model in Step 3 (e.g., hyperparameters, bootstrap sampling if using random forests, etc.).
-Once multiple imputed datasets have been created, inference can be done regarding a quantity of interest (e.g., mean of a variable, correlation coefficient between two variables).
-So-called Rubin’s rules provide statistical foundation for such inference.
+Once multiple imputed datasets have been created, inference can be done regarding a quantity of interest (e.g., the mean of a variable, the correlation coefficient between two variables).
+The so-called Rubin’s rules provide a statistical foundation for such inference.
 For a scheme that illustrates how MICE works, see Figure 1.
-For more information, see, e.g., Section 5.1 in Drechsler, (2011).
+For more information, see, e.g., Section 5.1 in Drechsler (2011).
 As mentioned earlier, the MICE framework can be used to generate synthetic datasets.
-For simplicity, we assume that the training data has no missing values, otherwise, a 2-step approach is needed, where before the generation (in the second step) in the first step, missing values are imputed.
+For simplicity, we assume that the training data has no missing values; otherwise, a 2-step approach is needed, where missing values are imputed in the first step before generation (second step).
 The idea is to iteratively select parts of the original data and then replace them with synthetic values drawn from the posterior distribution learned by MICE.
@@
 One way is to choose a full column in each iteration, cf.
-Volker and Vink, (2021).
+Volker and Vink (2021).
 This case has two peculiarities.
 First, the imputation model is trained on the observed data, which is then replaced by the model prediction.
-Second, the MICE algorithm converges immediately and each draw will be a direct draw from the posterior distribution, because a so-called monotone missingness pattern is satisfied for column-wise replacement/imputation (see Section 3.1.2 in Drechsler, (2011)).
-Another approach, which in general does not satisfy the monotone missingness pattern, is to randomly select in each amputation-imputation step cells whose values must be replaced.
-The values in these cells are deleted, interpreted as missing, and the MICE algorithm is executed, requiring more than one iteration to converge.
-In our case studies, we test several ways of iteratively amputating data and then applying MICE to impute it.
-One of them is inspired by the tabulator approach of Neves et al., (2022), where the authors suggest to amputate data in chunks of approximately the same size and to impute deleted data using a GAIN.
+Second, the MICE algorithm converges immediately, and each draw will be a direct draw from the posterior distribution, because a so-called monotone missingness pattern is satisfied for column-wise replacement/imputation (see Section 3.1.2 in Drechsler (2011)).
+Another approach, which in general does not satisfy the monotone missingness pattern, is at each amputation-imputation step to randomly select cells whose values must be replaced.
+The values in these cells are deleted, and treated as missing, and the MICE algorithm is executed, requiring more than one iteration to converge.
+In our case studies, we test several ways to iteratively set a portion of data to missing and then apply MICE to impute.
+One of them is inspired by the tabulator approach of Neves et al.
+(2022), where the authors suggest setting data to missing in chunks of approximately equal size and to impute missing data using a GAIN.
 Instead of a GAIN, we will use MICE as the imputation engine.
 Similarly to the MICE-based imputation, there are Rubin rules for analyzing (e.g., computing a point estimate, an estimate for the variance) unknown statistical quantities from fully synthetic datasets, cf.
-Reiter, (2003) and Section 6.1 in Drechsler, (2011).
+Reiter (2003) and Section 6.1 in Drechsler (2011).
 In our case studies, we use random forests (RFs) as the imputation model.
 This imputation model and many others are implemented in the R-package mice, cf.
-van Buuren and Groothuis-Oudshoorn, (2011).
-RFs were introduced in Breiman, (2001).
+van Buuren and Groothuis-Oudshoorn (2011).
+RFs were introduced in Breiman (2001).
 It is a non-parametric model that can capture complex relationships among variables without the need to specify a particular regression model.
@@
 Another important advantage of RFs is that they can often be used out-of-the-box and can be used rather easily to handle heterogeneous data.
-For more information on RFs in actuarial context, see, e.g., Chapter 4 in Denuit and Trufin, (2020) or Chapter 6 in Wuthrich et al., (2025).
+For more information on RFs in actuarial context, see, e.g., Chapter 4 in Denuit and Trufin (2020) or Chapter 6 in Wuthrich et al.
+(2025).
 3.2 Conditional Tabular GAN
-In this subsection, which is based on Kuo, (2020) and Sections $2$ and $3$ of Cote et al., (2020), we describe how a GAN works and explain how a CTGAN extends it.
+In this subsection, which is based on Kuo (2020) and Sections $2$ and $3$ of Côté et al. (2025), we describe how a GAN works and explain how a CTGAN extends it.
 A GAN consists of two neural networks (NNs): a generator $G$ and a discriminator $D$.
@@
 To improve the convergence of optimization routines for (3.1), researchers have proposed modifications to (3.1) such as using Wasserstein-$1$ distance between the true distribution $F_{\bm{X}}$ and the synthetic distribution $F_{G}$, enforcing additional constraints, adding penalties, etc.
-Proposed in Xu et al., (2019), CTGAN builds upon the GAN framework by incorporating several innovations to address the challenges of generating synthetic tabular data: mode-specific normalization, conditional generation of data, and handling of mixed data.
-Next we summarize an exemplary architecture of CTGAN, as in Cote et al., (2020).
-As per Kuo, (2020), a CTGAN generates individual records by randomly selecting a variable (e.g., fuel type) and then choosing a value (e.g., diesel) based on the actual data frequency.
+Proposed in Xu et al.
+(2019), CTGAN builds upon the GAN framework by incorporating several innovations to address the challenges of generating synthetic tabular data: mode-specific normalization, conditional generation of data, and handling of mixed data.
+Next we summarize an exemplary architecture of CTGAN, as in Côté et al.
+(2025).
+As per Kuo (2020), a CTGAN generates individual records by randomly selecting a variable (e.g., fuel type) and then choosing a value (e.g., diesel) based on the actual data frequency.
 The algorithm finds a matching training data row for this value and generates the remaining variables conditioned on the initial selection.
 Both the generated and true records are evaluated by the discriminator.
-Figure 3 in Cote et al., (2020) schematically illustrates this process and Figure 2 in Cote et al., (2020) visually exemplifies the architecture of the generator and discriminator.
+Figure 3 in Côté et al.
+(2025) schematically illustrates this process and Figure 2 in Côté et al.
+(2025) visually exemplifies the architecture of the generator and discriminator.
 The discriminator (blue) and the generator (orange) each have two fully-connected layers to capture relationships between columns.
-The generator, which supports categorical variables, employs skip connections; see Xu et al., (2019) for extensions to continuous variables.
-CTGAN also uses the PacGAN framework in the discriminator to mitigate a so-called mode collapse.
+The generator, which supports categorical variables, employs skip connections; see Xu et al.
+(2019) for extensions to continuous variables.
+CTGAN also uses the Packed Generative Adversarial Network (PacGAN) framework in the discriminator to mitigate a so-called mode collapse.
 This peculiarity means that the model focuses on a small subset of the real data and, thus, has a bias towards the generation of samples that are not sufficiently diverse.
-Although a CTGAN tends to be substantially better than a GAN for generating high-cardinality categorical variables, this type of variables can be still challenging for a CTGAN, depending on the characteristics of the real data such as the degree of cardinality and the frequency of each category.
+Although a CTGAN tends to be substantially better than a GAN for generating high-cardinality categorical variables, this type of variables can still be challenging for a CTGAN, depending on the characteristics of the real data, such as the degree of cardinality and the frequency of each category.
 Aiming to improve the performance of a CTGAN for high-cardinality categorical variables, we try encoding them via non-linear autoencoders (AEs) before training a CTGAN.
@@
 So principal components analysis approach can be interpreted as a linear AE, i.e., the activation functions of both NNs are linear.
-Hinton and Salakhutdinov, (2006) demonstrated the power of non-linear AEs for learning better, lower-dimensional representations than linear AEs.
+Hinton and Salakhutdinov (2006) demonstrated the power of non-linear AEs for learning better, lower-dimensional representations than linear AEs.
 In our paper, we use two different types of AEs due to their different application – deterministic AEs and variational AEs.
 Our application of deterministic AEs is the transformation of high-cardinality categorical variables to low-dimensional numerical vectors (instead of the one-hot encoding) before fitting them to GAN-based models for data generation.
-The benefits of using such AEs for training NNs in insurance predictive modeling has been studied in Delong and Kozak, (2023), which is our main reference for this application of non-linear AEs.
+The benefits of using such AEs for training NNs in insurance predictive modeling have been studied in Delong and Kozak (2023), which is our main reference for this application of non-linear AEs.
 Consider a vector of features $\bm{a}\in\mathbb{R}^{p}$, $p\in\mathbb{N}$.
@@
 Afterwards, this representation $\overline{\bm{x}}_{j}^{cat}=\varphi\left({\bm{x}}_{j}^{cat}\right)$, instead of ${\bm{x}}_{j}^{cat}$, is used as input for CTGAN to generate data.
-The approach of using a separate AE for each categorical variable is called Separate AEs in Delong and Kozak, (2023).
-We use the same architecture of separate AEs, as proposed by Delong and Kozak, (2023).
+The approach of using a separate AE for each categorical variable is called Separate AEs in Delong and Kozak (2023).
+We use the same architecture of separate AEs, as proposed by Delong and Kozak (2023).
 In particular, the reconstruction function $\pi$ is an NN with one hidden layer.
@@
 The approach described above is applied to all categorical features in the dataset.
-An example of the architecture of a neural network used in this paper to build the autoencoder of type Separate AEs for categorical features (with 2 and 3 categories) is presented in Figure 2 in Delong and Kozak, (2023).
+An example of the architecture of a neural network used in this paper to build the autoencoder of type Separate AEs for categorical features (with 2 and 3 categories) is presented in Figure 2 in Delong and Kozak (2023).
 See their Section 3.2 for further details on Separate AEs.
 Instead of using a separate AE for each categorical variable, we could have one joint AE for all categorical variables simultaneously.
-This approach is called Joint AE in Delong and Kozak, (2023).
-According to Figure 4 there, for a large number of epochs (e.g., $100$ for freMTPL2freq data) and a sufficiently high dimension of the representation of all categorical features (e.g., $15$ for freMTPL2freq data), the reconstruction error power measured with the cosine similarity is very similar for Separate AEs and Joint AE.
+This approach is called Joint AE in Delong and Kozak (2023).
+According to Figure 4 there, for a large number of epochs (e.g., $100$ for the freMTPL2freq data) and a sufficiently high dimension of the representation of all categorical features (e.g., $15$ for the freMTPL2freq data), the reconstruction error power measured with the cosine similarity is very similar for Separate AEs and Joint AE.
 The second type of AEs used in our paper is variational autoencoders (VAEs).
-They were introduced by Kingma and Welling, (2013) and can be seen as a Bayesian extension to deterministic AEs.
+They were introduced by Kingma and Welling (2013) and can be seen as a Bayesian extension to deterministic AEs.
 VAEs learn a probabilistic representation of data points in a latent space, which is why these models can be used for synthetic data generation in addition to representation learning.
-In our case studies, we also consider the VAE-based approach proposed by Jamotton and Hainaut, (2024), who also tested their methodology on a data set with about $62000$ observations.
+In our case studies, we also consider the VAE-based approach proposed by Jamotton and Hainaut (2024), who also tested their methodology on a dataset with about $62000$ observations.
 For an overview of VAEs and details about their methodology, we refer the reader to their article.
 4 Case studies
-This section is dedicated to our comparative study, which is based on freMTPL2freq dataset.
+This section is dedicated to our comparative study, which is based on the freMTPL2freq dataset.
 In Subsection 4.1, we provide the details about our method training pipelines, the mechanics of data augmentation, and the evaluation pipelines for generation methods.
@@
 This aspect is significant in practice since, in our experience from the industry, actuaries seldom have sufficient time to implement and fine-tune their specific algorithms and methods.
-Since we want to identify best “off-the-shelf” data generation approach, we do not do any pre-processing of variables when preparing data in this section.
+Since we want to identify the best “off-the-shelf” data generation approach, we do not do any pre-processing of variables when preparing data in this section.
 Any processing is done by the generation method implementation.
 We execute our case study on the high-performance computation cluster at the University of Tartu and can provide our code upon request.
-In our experiments, we use freMTPL2freq dataset to give us a baseline of what a realistic dataset might look like.
+In our experiments, we use the freMTPL2freq dataset to give us a baseline of what a realistic dataset might look like.
 It has $678013$ observations with $9$ explanatory variables ($5$ categorical and $4$ numeric), $1$ exposure variable and $1$ response variable (the number of claims).
-This dataset has been used in multiple actuarial data science articles, including Kuo, (2020), where it is used for training a CTGAN to generate synthetic data.
+This dataset has been used in multiple actuarial data science articles, including Kuo (2020), where it is used for training a CTGAN to generate synthetic data.
+The freMTPL2freq dataset is the de facto benchmark in modern actuarial data science, and introducing non-standard datasets would reduce comparability with existing literature.
 Although we have access to the original values of the observed response variable $y_{i}$ in the freMTPL2freq dataset, we do not know the true relation between the observed value $y_{i}$ and the corresponding observed covariates $\bm{x}_{i}$.
@@
 In the second formula we use the previously defined ${f}_{lin}$ and additionally assume a set of interactions among covariates, namely ${f}_{int}(\bm{x})={f}_{lin}(\bm{x})+\sum_{\{j,k\}\in\mathcal{I}}\beta_{j;k}x_{j}x_{k}$, where the interaction part is defined as in Equation (2.3) and $\mathcal{I}$ is the set of index pairs for the interactions of interest.
-Although we use double indices for notational simplicity, one can construct a simple mapping from two indices to single indices.
+Although we use double indices for notational clarity, one can construct a simple mapping from double indices to single indices.
 The values ${f}_{lin}(\bm{x}_{i})$ and ${f}_{int}(\bm{x}_{i})$ are then used as the means of the respective Poisson distributions to simulate the number of claims $y_{i}$, $i\in\left\{1,\dots,n_{ROD}\right\}$.
 So, in our experiments, we have two datasets, each of which contains simulated number claims for which we know their “true” structural dependence on original covariates.
-In the sequel, we refer to this artificially generated number of claims as the response variable and use the $90\%-10\%$ training-test split, i.e., the number of rows in the train set is $n_{RTR}=\lfloor n_{ROD}\cdot 0.9\rfloor$.
+In the sequel, we refer to this artificially generated number of claims as the response variable and use the $90\%/10\%$ training/test split, i.e., the number of rows in the train set is $n_{RTR}=\lfloor n_{ROD}\cdot 0.9\rfloor$.
 Let now $\mathcal{A}=\left\{A_{1},\dots,A_{n_{A}}\right\}$ be the set of algorithms (approaches) for generation of synthetic data, $n_{A}\in\mathbb{N}$ is the number of algorithms for synthetic data generation.
-For each approach $A$, we denote by $GM_{A}$ as the generation method for this algorithm.
+For each approach $A$, we denote the generation method for this algorithm by $GM_{A}$.
 For each algorithm, although the technical specifics of training can vary, the underlying pipeline we use is the same.
@@
 This requires additional steps in the training pipeline, which are outlined in Appendix A.
-Each generation method is used to generate $n_{E}$ (referred to as the number of experiments) synthetic datasets, where the number of rows for a single synthetic dataset was the same as the training data.
+Each generation method is used to generate $n_{E}$ (referred to as the number of experiments) synthetic datasets, where the number of rows for a single synthetic dataset is the same as in the training data.
 Denote a generated synthetic dataset as $\mathcal{S}^{A;k}$, where $k\in\{1,\dots,n_{E}\}$ and $A\in\mathcal{A}$.
@@
 The quantiles are used to ensure that the proportion of observations in each bin is the same or very close.
-Note that in the case when one value appears in several quantiles (e.g., one value is very prevalent in the data), then it has a bin all on its own.
+Note that if a value appears in several quantiles (e.g., it is very prevalent in the data), it has its own bin.
 For each bin, then, a binary variable (feature) is added, and the vector of these features is denoted by $\bm{x}_{j}^{cat}$.
@@
 For each dataset $\mathcal{S}^{A;k;\bm{s}}$, two predictive GLM models are derived.
-One model, where the formula used to build the model is the same as the corresponding true model formula (i.e.
-linear formula or interactions formula), and one which is derived using a bidirectional step-wise variable selection using $AIC$.
-Note that for interaction formula, the scope for the variable selection included variables explicitly defining the interactions between the different variables (since this was the easiest way to include interactions in the step-wise search).
+One model, where the formula used to build the model is the same as the corresponding true model formula (i.e., linear formula or interaction formula), and one which is derived using a bidirectional step-wise variable selection using $AIC$.
+Note that for the interaction formula, variable selection included explicitly defined columns for each of the interactions between the different variables, as this was the easiest way to include interactions in the step-wise search.
 The values related to the trained GLM models are denoted as follows:
@@
   • $\widehat{\bm{\beta}}^{A;k}=\left(\widehat{\beta}_{0}^{A;k},\widehat{\beta}_{1}^{A;k},\dots,\widehat{\beta}_{d}^{A;k}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained only on the synthetic dataset $\mathcal{S}^{A;k}$.
-  • $\widehat{\bm{\beta}}^{A;k;\bm{s}}=\left(\widehat{\beta}_{0}^{A;k;\bm{s}},\widehat{\beta}_{1}^{A;k;\bm{s}},\dots,\widehat{\beta}_{d}^{A;k;\bm{s}}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained on trained on the dataset specified by the data structure parameter $s$.
+  • $\widehat{\bm{\beta}}^{A;k;\bm{s}}=\left(\widehat{\beta}_{0}^{A;k;\bm{s}},\widehat{\beta}_{1}^{A;k;\bm{s}},\dots,\widehat{\beta}_{d}^{A;k;\bm{s}}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained on the dataset specified by the data structure parameter $s$.
   • $MSE\left(\widehat{\beta}_{j}^{\bullet}\right)$ – mean squared error of the estimated coefficient $\widehat{\beta}_{j}^{\bullet}$, where $\bullet$” represents any algorithm/data structure combination.
-As these values can be obtained for every trained GLM model, we use two metrics to assess the effect of synthetic data on estimated values of model coefficients.
+As these values can be obtained for every trained GLM model, we use two metrics to assess the effect of synthetic data on the estimated values of model coefficients.
 First, to measure the distance between the true GLM coefficients and their estimate based on the true structure of a GLM and a synthetic dataset generated by an approach $A$ and specified by the structure parameter $\bm{s}$, we use the following metric:
@@
 The first group of methods we test are GAN-based, namely:
-  • CTGAN (CTGAN) – implementation of Xu et al., (2019) available in the sdv library for Python;
+  • CTGAN (CTGAN) – implementation of Xu et al. (2019) available in the sdv library for Python;
   • CTGAN with AEs (CTGAN_WITH_AE) – a CTGAN enhanced by AEs for encoding-decoding categorical variables;
-  • Wasserstein GAN with gradient penalty (COTE_MC_WGAN_GP) – implementation of the method graciously provided by Cote et al., (2020);
-  • GAN-based tabulator (JCS_TABULATOR) – tabulator as proposed and implemented by Neves et al., (2022).
+  • Wasserstein GAN with gradient penalty (COTE_MC_WGAN_GP) – implementation of the method graciously provided by Côté et al. (2025);
+  • GAN-based tabulator (JCS_TABULATOR) – tabulator as proposed and implemented by Neves et al. (2022).
 The second group of synthetic-data generators we test are the imputation-amputation methods based on MICE.
-One of the considered approaches is tabulator proposed Neves et al., (2022), but we use MICE as the imputation engine instead of GAIN-based models suggested there.
-In our “MICE-tabulator”, we have $5$ sequential amputation-imputation iterations, each of each amputates randomly selected $20\%$ of cells and imputes them to generate synthetic data points.
+One of the considered approaches is the tabulator proposed by Neves et al.
+(2022), but we use MICE with RFs as the imputation engine instead of the GAIN-based models suggested there.
+In our MICE-based tabulator, we have $5$ sequential amputation-imputation iterations, each of which sets random $20\%$ of cells to missing and imputes them using MICE-RF to generate synthetic data points.
 The other two methods have a different amputation logic.
-First, we create a copy of the training data, where we remove $75\%$ of cell values.
-Second, we append this data with missing values to the original training data and apply MICE to impute the missing values.
-The data where missing cells were present, is treated as the generated data.
-In this way, we end up with generated data where 25% of the data comes from the original training dataset and 75% is synthetic.
-Hereinafter, we refer to this approach as the “MICE method”.
-The third approach extends the MICE method by the second iteration, in which the cells that were not amputated-imputed in the first iteration are amputated and imputed via MICE.
-This leads to a data set that is fully synthetic.
-In summary, the MICE-based methods we consider are
-  • MICE method (MICE) – the method described above and implemented using the R-package mice;
-  • MICE with all synthetic data (MICE_ALL_SYN) – the modification of the MICE method, where in an additional step the remaining $25\%$ of the original data is imputed via MICE;
-  • MICE tabulator (MICE_TABULATOR) – the tabulator setup from Neves et al., (2022) with $5$ iterations and MICE with RFs as the imputation engine.
-After testing the above-described groups of methods, we noticed that GAN-based approaches learned the univariate distribution of numeric variables worse than MICE-based methods and had challenges with high-cardinality categorical variables.
-Due to this finding, we decided to consider in our case studies two hybrid models.
-One of them combines a CTGAN and a MICE component, namely a CTGAN first generates synthetic data and then $75\%$ of synthetic numeric variable values and imputed amputated and imputed using MICE.
+First, we create a copy of the training data, where we set $75\%$ of cell values to missing.
+Second, we append this data with missing values to the original training data and apply MICE-RF to impute the missing values.
+The data with missing cells is treated as generated data.
+In this way, we end up with generated data where 25% of the data comes from the original training dataset, and 75% is synthetic.
+Although the resulting dataset is partially synthetic, the probability of disclosing confidential identifiable information or of having the same row in both the synthetic data and the original data should be negligible due to the completely-at-random way of creating missingness in the original data.
+Hereinafter, we refer to this approach as the “MICE partially synthetic method”.
+The third approach extends the MICE partially synthetic method by the second iteration, in which the cells that were not imputed in the first iteration (the remaining quarter of the original data) are set to missing and then imputed via MICE.
+This results in a fully synthetic dataset.
+This method is referred to as “MICE fully synthetic method”.
+For each of the approaches described in this paragraph, we use only one synthetic dataset for evaluating data fidelity and data utility.
+This is different from the original idea of MICE for statistical inference (see Subsection 3.1, Figure 1), where multiple different imputed datasets are created and then combined to study a statistical object with the help of Rubin’s rules.
+In summary, the MICE-based methods we explore and benchmark in our case study are:
+  • MICE partially synthetic method (MICE_PART_SYN) – the method described above and implemented using the R-package mice;
+  • MICE fully synthetic method (MICE_FULL_SYN) – the modification of the partially synthetic method, where in an additional step the remaining $25\%$ of the original data is imputed via MICE;
+  • MICE tabulator (MICE_TABULATOR) – the tabulator setup from Neves et al. (2022) with $5$ iterations and MICE with RFs as the imputation engine.
+  • MICE as per Volker-Vink (MICE_VV) – amputation-imputation procedure utilizing MICE columnwise as described in Volker and Vink (2021).
+Before we move on to other synthetic data generators in our case study, we clarify the different types of “iterations” in MICE-based methods.
+There are amputation-imputation iterations and there are MICE iterations.
+The former iteration type means that, first, a not yet imputed part of the training data is set to missing and, second, the respective missing values are imputed (using MICE-RF in our case study).
+A MICE iteration should be understood as a sequence of actions in Steps 2-5 in Figure 1 done one single time, i.e., once for each variable creating missingness, training an RF that uses other variables to predict the variable with missingness, and replacing introduced missing values by RF predictions.
+For example, in our case study, the MICE-tabulator approach has $5$ amputation-imputation iterations, the MICE partially synthetic method has $1$ amputation-imputation step, and the MICE fully synthetic method has $2$ amputation-imputation steps.
+In the MICE-based generation as per Volker and Vink (2021), there is no explicit creation of missingness before training imputation models.
+Instead, for each column, an imputation model is trained without randomly deleting observed values.
+Then each observed value is set to missing and then replaced by the respective prediction from the trained imputation model.
+The resulting imputed column is then used as the predictor for further imputation in the chain of columns.
+Since a single pass over the original dataset is performed only once to obtain a fully synthetic dataset, there are in total $n_{C}$ amputation-imputation iterations.
+After testing the above-described groups of methods, we noticed that GAN-based approaches learned the univariate distributions of numeric variables worse than MICE-based methods and had challenges with high-cardinality categorical variables.
+Due to this finding, we decided to consider two hybrid models.
+One of them combines a CTGAN and a MICE component, namely a CTGAN first generates synthetic data and then $75\%$ of the synthetic numeric variable values are set to missing and then imputed using MICE.
 Note that also for this CTGAN-MICE model, the data with removed cell values was combined with the original training data to impute the missing numerical column values.
@@
   • CTGAN with AEs and MICE (CTGAN_WITH_AE_MICE) – a CTGAN enhanced by AEs for categorical variables and MICE for numeric variables.
-Lastly, we also test the method proposed in Jamotton and Hainaut, (2024), where variational autoencoders (VAEs) are used for data generation.
-We use the implementation graciously shared with us by the authors of Jamotton and Hainaut, (2024).
+Lastly, we test the method proposed in Jamotton and Hainaut (2024), which uses variational autoencoders (VAEs) for data generation.
+We use the implementation graciously shared with us by the authors of Jamotton and Hainaut (2024).
 The monospaced shorthand for this method is VAE_JAMOTTON.
-As described in Subsection 4.1, artificial claim counts are generated for this case study such a way that we know the true relationship between the response variable and the descriptive variables.
+As described in Subsection 4.1, artificial claim counts are generated for this case study in such a way that we know the true relationship between the response variable and the descriptive variables.
 The following formula for a linear relationship is used:
@@
 $-0.3\cdot\mathbbm{1}_{\left(\text{AREA}\in\text{\{F\}}\right)}\cdot\mathbbm{1}_{\left(\text{VEHICLE\_BRAND}\in\text{\{B1,B2,B12\}}\right)}.$
-The claim counts simulated using (4.3) and (4.2) lead to the average claim frequency of $5.37\%$ and $6.04\%$ respectively.
-These numbers are close to the average claim frequencies that we have observed in the industry when modeling MTPL claims.
+The claim counts simulated using (4.3) and (4.2) lead to the average claim frequency of $5.37\%$ and $6.04\%$, respectively. These numbers are close to the values observed in practice.
+We acknowledge that the chosen linear and interaction formulas follow the setup of generalized linear models, which is the core basis of insurance ratemaking in practice.
+The topic of more complex dependence structures is left as an avenue for further research.
 4.3 Case study results
@@
 We do not claim that one particular method is best suited for synthetic data generation, but we assess the strengths and weaknesses of different methods.
-With our case study, we evaluated a total of $10$ different approaches to synthetic data generation.
-However, we display the results only for $9$ of them, as the tested tabulator by Neves et al., (2022) did not generate any number of claims, and, therefore, it could not be fully evaluated.
+With our case study, we evaluated a total of $11$ different approaches to synthetic data generation.
+However, we display the results only for $10$ of them, as the tested tabulator by Neves et al.
+(2022) did not generate any number of claims, and, therefore, it could not be fully evaluated.
 In total, we have two sets of results, one for the case linear formula (4.3), and one for the case where the interaction formula (4.2) is used.
@@
 Data-specific metrics are MAE/MAPE for numeric, categorical and pair-wise combinations of columns and MAE/MAPE for correlation structure within numeric columns of the generated data.
-Lastly, our subjective rank of the method’s ease of use is also provided.
-Note that every synthetic dataset is uniquely generated 5 times ($n_{E}=5$) and all of the following results are presented as simple averages across all generated datasets for a given method (except for $M_{1}$ and $M_{2}$ metrics, as these are already averaged across $n_{E}$).
-For data-specific metrics, every column is taken as equally important, so the shown metric is an average taken over all columns or column combinations.
+Lastly, our subjective ranking of the method’s ease of use is also provided.
+Note that every synthetic dataset is uniquely generated 5 times ($n_{E}=5$) and all of the following results are presented as simple averages across all generated datasets for a given method (except for $M_{1}$ and $M_{2}$ metrics, since these incorporate $n_{E}$).
+For data-specific metrics, each column is taken as equally important, so the metrics reflect the average of all columns or combinations.
 The final results for the evaluation of methods are presented in Tables 1 and 2.
 Each cell shows the rank of the method with respect to the metric shown in the column.
-Additionally, the table is color matched based on the rank of the method, and in the case of ties, with respect to metric, all cells are given the same highest rank available.
+Additionally, the table is color-matched based on the rank of the method, and in the case of ties, with respect to the metric, all cells are given the same highest rank available.
 Note that in all cases, the ”training” method is shown as the reference value for different metrics.
+Before looking at the results, it is important to comment that in our case study all generators were trained using the University of Tartu high-performance computation cluster (University of Tartu (2018)).
+Because this cluster does not inherently have hardware stability, we were unable to directly compare fitting and generation times.
+However, we observed that MICE-based and CTGAN-based methods usually took the least time (around 3 hours for model fitting and data generation), while more complex custom implementations, such as COTE_MC_WGAN_GP, took significantly longer for model fitting and generation (about 10 hours in total).
 Generation method M1 M2
@@
 1
-(3,15668)
-1
-1
-1
-(0,31917)
-1
-(0,23744)
 1
@@
 1
-MICE
+1
+1
+1
+1
+MICE_PART_SYN
 2
-(33,93479)
 2
-(33,93813)
+1
+5
+2
+2
+2
+3
+4
+2
+5
+3
+3
+3
+2
+MICE_FULL_SYN
+3
+3
 1
 4
-(1,2)
-2
-(0,32027)
-2
-(0,2376)
-2
-(0,00061)
-2
-(0,33907)
-4
-(0,00945)
-2
-(0,27654)
-5
-(0,00185)
-3
-(0,02183)
-3
-(0,0011)
-2
-(0,27231)
-2
-CTGAN
-3
-(41,52358)
-3
-(41,03633)
-1
 3
 3
-(0,32039)
 3
-(0,23761)
+5
+5
 3
-(0,00072)
+6
 4
-(0,44582)
+4
+6
+4
+MICE_VV
 7
-(0,01192)
-3
-(0,33968)
-9
-(0,00678)
-9
-(0,11463)
-4
-(0,00156)
-5
-(0,37186)
-3
-MICE_ALL_SYNTHETIC
-4
-(114,40835)
-5
-(112,49103)
-9
-2
-(0,8)
-4
-(0,32396)
-4
-(0,23812)
-4
-(0,00107)
-5
-(0,4497)
-5
-(0,01031)
-4
-(0,35079)
-6
-(0,00232)
-4
-(0,02677)
-2
-(0,000995)
-4
-(0,36295)
-3
-MICE_TABULATOR
-5
-(111,53667)
-4
-(113,48165)
-10
-(5,6)
 8
 6
-(0,32592)
+2
+5
+7
+5
+2
+10
+8
+7
+5
+5
+2
+2
+MICE_TABULATOR
+4
+5
+10
+3
+4
+4
+4
 6
-(0,23834)
+9
+6
+2
+2
+2
+5
+4
+CTGAN
+11
+11
+1
 7
-(0,00253)
+10
+10
+9
+9
+7
+10
+10
+10
+8
+9
+4
+CTGAN_MICE
+5
+4
+11
+9
+7
 6
-(3,96253)
+8
+7
+2
+4
+10
+10
+6
+7
+7
+CTGAN_WITH_AE
+8
+6
+6
 9
-(0,0249)
+11
+11
 7
-(0,67136)
-2
-(0,00081)
-2
-(0,01022)
-5
-(0,00707)
+10
+8
+11
+3
 6
-(18,16422)
+9
+10
+7
+CTGAN_WITH_AE_MICE
+10
+9
+6
+7
+9
+9
+6
+8
 3
+7
+3
+6
+7
+8
+9
 VAE_JAMOTTON
 6
-(129,11846)
 7
-(126,94602)
+9
+5
+6
+5
+10
+4
+11
+9
 8
-(6,6)
-4
-(1,2)
-5
-(0,32545)
-5
-(0,23829)
-9
-(0,00303)
-3
-(0,40646)
-10
-(0,03651)
-9
-(1,07215)
-7
-(0,00429)
-7
-(0,0809)
-9
-(0,02964)
-3
-(0,32234)
-9
-COTE_MC_WGAN_GP
-8
-(155,80504)
-9
-(152,74518)
-1
-8
-7
-(0,33124)
-7
-(0,23896)
-10
-(0,00412)
-10
-(82,61544)
-6
-(0,01052)
-6
-(0,48478)
-8
-(0,00444)
-8
-(0,08287)
-10
-(0,04028)
-10
-(445,05061)
-10
-CTGAN_MICE
-10
-(393,44698)
-10
-(383,62816)
-1
-6
-(1,4)
-9
-(0,34255)
-9
-(0,23977)
-8
-(0,00292)
-8
-(8,9032)
-2
-(0,00543)
-5
-(0,44177)
-9
-(0,00678)
-9
-(0,11463)
-7
-(0,0147)
-8
-(42,48769)
-6
-CTGAN_WITH_AE_MICE
-9
-(143,03862)
-8
-(153,73562)
-6
-(6,8)
-6
-(1,4)
-8
-(0,33266)
-8
-(0,23937)
-5
-(0,00199)
-7
-(8,25317)
-3
-(0,00569)
-8
-(0,71934)
-3
-(0,0016)
-5
-(0,05566)
-6
-(0,00974)
-7
-(38,72062)
-8
-CTGAN_WITH_AE
-7
-(125,10637)
-6
-(136,08475)
-6
-(6,8)
 8
 10
-(0,34588)
+4
 10
-(0,24254)
+COTE_MC_WGAN_GP
+9
+10
+1
+9
+8
+8
+11
+11
 6
-(0,00244)
+5
 9
-(17,68936)
-8
-(0,01736)
-10
-(1,74036)
-3
-(0,0016)
-5
-(0,05566)
-8
-(0,02108)
 9
-(83,98836)
-6
+11
+11
+11
 Table 1: Summary of evaluation results across all tested algorithms when the variable dependence is based on linear formula (4.3).
@@
 The differences between training data and synthetic data can be quite large for example for metric $M_{1}$ or $M_{2}$.
-However, when it comes to the overall accuracy of the model, the difference in coefficients does not result in a very significant difference in the accuracy of the predictions.
-We can see that the best-performing synthetic data generation methods are MICE and CTGAN.
-They show almost the same accuracy in terms of goodness-of-fit metrics, but MICE is better when it comes to the accuracy of model coefficients.
-The worst performing methods are CTGAN_MICE, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE.
-All of these methods not only show worse accuracy in model coefficients but also the model predictions are worse compared to other methods.
+However, when it comes to overall model accuracy, the difference in coefficients does not translate into a significant difference in prediction accuracy.
+We can see that the best-performing synthetic data generation methods are MICE_PART_SYN and MICE_FULL_SYN.
+They show almost the same accuracy in terms of goodness-of-fit metrics, but MICE_PART_SYN is better when it comes to the accuracy of model coefficients.
+The worst performing methods are CTGAN, CTGAN_WITH_AE_MICE, COTE_MC_WGAN_GP and CTGAN_WITH_AE.
+These methods have higher model coefficient metric values and worse accuracy in model predictions.
 Almost all methods show consistency in the correct terms using the variable selection method.
-However, for all methods the variable selection included additional variables not present in the true model.
-Here, the worst performing method was MICE_TABULATOR, where automatic variable selection on average failed to include 1.4 correct variables and included, on average, 2 additional incorrect variables (the worst possible result) to the model.
-The best variable selection results were achieved by models trained on data using the CTGAN method, always selecting the correct variables and including only 1 additional incorrect variable on average.
-Similar observations can be seen when it comes to data-specific metrics.
-Here again, MICE performs the best across all data metrics.
+However, for all methods, the variable selection included additional variables not present in the true model.
+Here, the worst-performing method was CTGAN_MICE, where automatic variable selection failed to include 1.4 correct variables, on average, and included, on average, 2 additional incorrect variables (the worst possible result).
+The best variable selection results were achieved by models trained on data using the MICE_PART_SYN and MICE_FULL_SYN methods, which always selected the correct variables and included only 1 additional incorrect variable on average.
+Notably, MICE_VV included, on average, only $0.4$ incorrect variables.
+Similar observations can be seen for data-specific metrics.
+Here again, MICE_PART_SYN overall performs the best across all data metrics.
 From Table 1, we conclude that different methods have different niches.
 For example, it seems MICE_TABULATOR is much better at categorical column generation, followed closely by CTGAN_WITH_AE and CTGAN_WITH_AE_MICE (since they share categorical variables).
-This possibly shows that the usage of autoencoders improves the generation of categorical columns.
-Additionally, we can observe that the addition of MICE to CTGAN-based methods might not be very consistent.
-MICE helps in terms of numerical column accuracy for CTGAN_WITH_AE_MICE, but makes it worse for CTGAN method.
-WGAN by Cote et al., (2020) shows overall the worst performance when it comes to data-specific metrics.
-As for to the ease of use, MICE is one of the easiest methods to apply, since the implementation in the R package mice is straightforward and requires no additional implementation or data preparation besides the choice of parameters for the function.
-CTGAN-based methods seem to require more tinkering and also a change of environment to Python since the implementation of the CTGAN interface in R (through the package ctgan) in our experience is neither robust nor intuitive.
-Addition of autoencoders, although simple in idea, can be quite hard to implement since this requires additional steps with respect to data preparation, setting up and fitting the additional neural networks and finally clustering the generated vectors to get back to the correct categorical labels.
-The most complex implementations were those provided by Jamotton and Hainaut, (2024) and Cote et al., (2020) since they required the most additional work to make it work properly in our case study.
+This suggests that the use of autoencoders improves the generation of categorical columns.
+Additionally, we observe that adding MICE to CTGAN-based methods slightly improves the generation of numeric columns, as both CTGAN_MICE and CTGAN_WITH_AE_MICE perform better than their non-augmented counterparts.
+WGAN by Côté et al.
+(2025) shows the worst overall performance on data-specific metrics.
+We also qualitatively assess the usability of each tested generator based on three criteria: intuitiveness (conceptual clarity), operational complexity (training and usage pipelines), and deployment overhead (environment setup, debugging, and code adaptation).
+We find MICE-based methods to be the most accessible, primarily due to their streamlined implementation in the R package mice, which requires minimal setup beyond standard installation.
+Notably, mice outputs by default synthetic data without exporting the trained imputation models.
+In contrast, CTGAN-based methods require additional setup time and more extensive data preprocessing (e.g., matrix fragmentation, type casting).
+For practical reasons, these experiments were conducted in Python, as this environment was more compatible with our computational workflow than the available R implementation.
+Incorporating AEs, while conceptually straightforward, increases implementation complexity due to the need for supplementary data preparation, neural network specification, and post-hoc clustering to recover categorical labels.
+Finally, the custom implementations kindly provided by the authors of Jamotton and Hainaut (2024) and Côté et al.
+(2025) represent highly specialized, advanced architectures.
+Consequently, integrating these custom-built models into a standardized, comparative computational framework naturally necessitates a higher degree of code adaptation and environment configuration.
 Generation method M1 M2
@@
 1
-(0,86047)
-1
-1
-1
-1
-(0,3394)
-1
-(0,25286)
 1
@@
 1
-MICE
+1
+1
+1
+1
+1
+MICE_PART_SYN
 2
-(17,6173)
 2
-(13,62087)
 3
-(15,4)
-6
-(1,8)
+7
 1
 2
-(0,34127)
 2
-(0,25318)
 2
-(0,00061)
+4
+4
+2
 3
-(0,34177)
+3
+3
 4
-(0,00889)
 2
-(0,30741)
+MICE_FULL_SYN
 3
-(0,002)
 3
-(0,02169)
+2
+5
+6
 3
-(0,00139)
 3
-(0,29525)
+3
+5
+5
+4
+4
+4
+4
+6
+4
+MICE_VV
+9
+9
+6
+3
+6
+6
+6
+5
 2
+11
+8
+5
+5
+5
+2
+2
+MICE_TABULATOR
+8
+8
+8
+8
+11
+5
+5
+4
+6
+10
+6
+2
+2
+2
+5
+4
 CTGAN
+11
+10
+4
+5
 3
-(21,04755)
-3
-(16,69526)
-2
-(15,6)
+11
+10
+10
+8
+7
+11
+10
+10
+10
+8
 4
-(1,6)
-6
-(0,4)
-3
-(0,34166)
-3
-(0,25323)
-3
-(0,00073)
-4
-(0,45389)
-7
-(0,01121)
-3
-(0,37872)
-9
-(0,$01$)
-9
-(0,09023)
-4
-(0,00175)
-5
-(0,41926)
-3
-MICE_ALL_SYNTHETIC
-8
-(55,7844)
-8
-(49,90837)
-7
-(9,4)
-7
-10
-(1,2)
-5
-(0,34909)
-5
-(0,25434)
-4
-(0,00111)
-5
-(0,4616)
-5
-(0,00967)
-5
-(0,38692)
-4
-(0,00242)
-4
-(0,02565)
-2
-(0,00102)
-4
-(0,3482)
-3
-MICE_TABULATOR
-7
-(53,21122)
-6
-(49,87283)
-8
-(8,4)
+CTGAN_MICE
 7
 6
-(0,4)
+9
 8
-(0,35463)
+6
+9
+9
+9
+7
+2
+3
+10
+10
+7
+7
+7
+CTGAN_WITH_AE
+4
+4
+7
+4
+3
+4
+4
 8
-(0,25515)
+10
 8
-(0,00243)
+10
+8
 6
-(2,93129)
+9
 10
-(0,02481)
 7
-(0,49791)
-2
-(0,00071)
-2
-(0,00955)
+CTGAN_WITH_AE_MICE
+5
+5
+10
+8
 6
-(0,005)
+7
+7
 6
-(13,22874)
+9
 3
+5
+8
+6
+6
+9
+9
 VAE_JAMOTTON
 6
-(55,09532)
 7
-(49,58059)
-10
-(7,8)
+11
 2
-(0,4)
 6
-(0,4)
-7
-(0,35376)
-7
-(0,25501)
-6
-(0,00203)
-2
-(0,31078)
-9
-(0,02237)
 8
-(0,95891)
-6
-(0,00523)
-7
-(0,07658)
-7
-(0,00784)
-2
-(0,21804)
-9
-COTE_MC_WGAN_GP
-9
-(101,62427)
-10
-(95,74713)
-5
-(11,8)
+8
 7
 3
-(0,2)
-9
-(0,3585)
-10
-(0,25585)
-10
-(0,00447)
-10
-(49,09032)
-6
-(0,0105)
-9
-(1,30688)
-5
-(0,00448)
-8
-(0,08906)
-10
-(0,02414)
-10
-(312,41321)
-10
-CTGAN_MICE
-10
-(97,45528)
-9
-(97,22979)
-4
-4
-(1,6)
-3
-(0,2)
-10
-(0,37112)
-9
-(0,25564)
-9
-(0,00278)
-7
-(6,29387)
-2
-(0,00479)
-4
-(0,38592)
-9
-(0,$01$)
-9
-(0,09023)
-9
-(0,01193)
-7
-(30,54492)
-6
-CTGAN_WITH_AE_MICE
-5
-(50,98831)
-5
-(47,65917)
 9
 7
+7
+8
+8
+3
+10
+COTE_MC_WGAN_GP
+10
+11
+5
+8
+3
+10
+11
+11
+11
 6
-(0,4)
+9
 6
-(0,35251)
-6
-(0,25483)
-5
-(0,00194)
-8
-(8,69745)
-3
-(0,00555)
-6
-(0,47747)
-7
-(0,00529)
-5
-(0,06383)
-5
-(0,00387)
-8
-(45,25532)
-8
-CTGAN_WITH_AE
-4
-(46,39565)
-4
-(41,8)
-6
-3
-(1,2)
-3
-(0,2)
-4
-(0,34617)
-4
-(0,2539)
-7
-(0,00207)
 9
-(19,89066)
-8
-(0,02117)
-10
-(1,79309)
-7
-(0,00529)
-5
-(0,06383)
-8
-(0,00852)
-9
-(103,25122)
-6
+11
+11
+11
 Table 2: Summary of evaluation results across all tested algorithms when the variable dependence is based on interaction formula (4.2).
@@
 In other words, this indicates that the generator fidelity depends on the underlying data dependence structure.
-From model-specific results, we see that coefficient distance metrics $M_{1}$ and $M_{2}$ have much smaller values compared to the linear case.
-The best methods based on coefficient difference metrics are again MICE and CTGAN, and the worst method seems to be COTE_MC_WGAN_GP and CTGAN_MICE.
+From the model-specific results, we see that the coefficient distance metrics $M_{1}$ and $M_{2}$ have much smaller values than in the linear case.
+The best methods based on coefficient difference metrics are again MICE_PART_SYN and MICE_FULL_SYN, and the worst method seems to be COTE_MC_WGAN_GP and CTGAN.
 Notably, CTGAN_WITH_AE and CTGAN_WITH_AE_MICE show a big uplift in rankings while additional MICE-based methods lose ranks compared to the linear case.
-This possibly shows that MICE is less suited for data structures with inherent dependence, but overall performance still seems to be the best across the tested approaches.
-Model goodness-of-fit metrics reflect very similar patterns as model coefficient metrics.
-However, MICE_ALL_SYNTHETIC shows a higher rank with respect to the goodness-of-fit metrics.
-The Poission deviance for it is notably smaller (0.00554) than that for MICE_TABULATOR.
-The variable selection shows, however, much more varied results.
-Again, MICE and CTGAN methods show best results, where variable selection, on average, selected 15.4 and 15.6 correct terms.
-Beyond these two methods, the variable selection in the next best method selected, on average, 12 correct terms.
-Variable selection for VAE_JAMOTTON got the worst result, on average, selecting only 7.8 correct terms to the model.
+This suggests that MICE is less suited to data structures with inherent dependence, but overall performance still seems best across the tested approaches.
+Model goodness-of-fit metrics yield slightly different results than model coefficient metrics.
+MICE-based methods lead in terms of these metrics, while GAN and VAE-based methods show worse results.
+Notably, CTGAN_WITH_AE shows much better results when compared to other GAN-based generation methods.
+The variable selection, however, shows much more varied results.
+Again, MICE_PART_SYN and MICE_FULL_SYN methods show the best results, where variable selection, on average, selected 15.4 and 15.6 correct terms.
+Beyond these two methods, the next-best method selected, on average, 12 correct terms.
+Variable selection for VAE_JAMOTTON got the worst result, on average, selecting only 7.8 correct terms for the model.
 We can also see that for almost all methods, variable selection added additional incorrect terms, where only VAE_JAMOTTON picked on average 0.4 additional incorrect terms.
-These combined facts indicate that using VAE_JAMOTTON results in a small/lean model, which is good for the linear case, but worse for the interaction case.
-From the data-specific results, we can observe almost the same rankings as we did for the linear case.
-This is expected since we did not introduce any additional inter-variable relationships besides the interaction relationship between response and the two covariates.
-The biggest difference is the fact that autoencoders did not improve the MAE/MAPE of categorical columns as significantly for CTGAN method as they had done in the linear case study.
-MICE is again clearly the leader when it comes to data-specific metrics, achieving top ranks in all metrics.
-However, analyzing synthetic data on a more granular level, we noticed that MICE-based methods can generate numeric values that ignore business constraints, e.g., generating observations with a driver age below $18$.
+These combined facts indicate that using VAE_JAMOTTON results in a small/lean model, which is not desirable in this case.
+From the data-specific results, we observe rankings that are almost identical to those for the linear case.
+This is expected, since we did not introduce any additional inter-variable relationships beyond the interaction between the response and the two covariates.
+Here again, the use of autoencoders improves MAE/MAPE for categorical columns; however, this improvement is not sufficient to beat the results of MICE-based methods.
+Again, MICE_TABULATOR shows very strong performance on categorical columns.
+Additionally, MICE_VV and MICE_TABULATOR indicate a loss of the original data correlation structure.
+Upon analyzing synthetic data on a more granular level, we noticed that MICE-based methods can generate numeric values that ignore business constraints, e.g., generating observations with a driver age below $18$.
 Imposing such constraints may be investigated in a follow-up paper.
+Complementing the results in Table 1 and Table 2, we analyzed the averaged (across all experiments) covariance matrices for numeric variables in the generated data.
+The resulting covariance matrices support the findings above, identifying MICE_PART_SYN and MICE_PART_SYN methods give, on average, the closest covariance matrices to the covariance matrix for the original numerical variables.
+However, it is important to note that GAN-based models performed better than MICE-based methods with respect to the variance of the density column.
+In particular, CTGAN maintained accurate variance, whereas the MICE-based methods underestimated this dispersion measure compared to the original data.
 Figure 2: $M_{1}$ metric results as a function of synthetic data proportion.
@@
 Having discussed data generation fidelity, we now focus on data augmentation.
-For this, we look at $M_{1}$ metric depending on the proportion of synthetic data added.
-We observed that as we increase the proportion of additional synthetic data compared to the original training data, the value for $M_{1}$ metric increases.
-The only case where we could observe a decrease in a metric was for coefficient distance metrics $M_{1}$ for the method MICE with an additional synthetic data proportion of $0.2$.
+For this, we examine how the value of the $M_{1}$ metric depends on the proportion of synthetic data added.
+We observe that as we increase the proportion of additional synthetic data compared to the original training data, the value for the $M_{1}$ metric increases.
+The only case in which we observed a decrease in a metric is for the metric $M_{1}$ for the method MICE_PART_SYN and a synthetic data proportion of $0.2$ in the interaction case .
 In this case, the augmented data achieved the $M_{1}$ metric value of $0.78717$ while the corresponding training data metric was $0.86047$.
-Since this is the only result where this happened, it is not enough to say that augmentation of the original freMTPL2freqdata might help when it comes to better accuracy of coefficients.
-The general relationship between the increase of additional synthetic data proportion and metric value can be seen in Figure 2.
-The shown figure is for metric $M_{1}$.
-As the synthetic data proportion increases, we can see a clear increasing trend for $M_{1}$ metric value for the given method.
-Based on the way $M_{1}$ metric is calculated and how GLM relies on the underlying data to compute the coefficient values, we estimate that if we keep on increasing the proportion of additional synthetic data, then we will asymptotically reach the metric values which we can compute for case where we use only synthetic data.
+This is the only result where this happened, it is not enough to say that augmentation of the original freMTPL2freq data might help when it comes to better accuracy of coefficients.
+The general relationship between the increase in the proportion of additional synthetic data and the metric $M_{1}$ value is shown in Figure 2.
+As the proportion of synthetic data increases, we observe a clear upward trend in the $M_{1}$ metric value for the given method.
+Based on the way the $M_{1}$ metric is calculated and how GLM relies on the underlying data to compute the coefficient values, we estimate that if we keep on increasing the proportion of additional synthetic data, then we will asymptotically reach the metric values which we can compute for the case where we use only synthetic data.
 These pure synthetic data results are shown on the right side of the plot.
@@
 Our paper demonstrates that the amputation-imputation framework based on MICE can be effectively used within an actuarial context.
-This framework is competitive for generating synthetic ratemaking data when compared to more sophisticated deep generative models such as CTGANs or VAEs.
+This framework is competitive with more sophisticated deep generative models, such as CTGANs or VAEs, for generating ratemaking data.
 In our case studies based on the open-source freMTPL2freq dataset, synthetic data generated using MICE with random forests as the imputation engine shows, on average, the best properties in terms of modeling efficacy and achieves high ranks in model accuracy and coefficient difference metrics.
-In general, several of the $9$ tested methods for synthetic data generation seem to have their niches for specific use cases.
+In general, several of the $10$ tested methods for synthetic data generation seem to have their niches for specific use cases.
 Having subjectively evaluated the ease of use for each tested method, we identify MICE as the most user-friendly approach, since it requires the least amount of additional work to get started, which is an important practical aspect for actuaries.
-Custom deep generative models like WGAN by Cote et al., (2020) or VAE by Jamotton and Hainaut, (2024) seem to require significantly more expertise and tuning to get good results.
+Custom deep generative models like WGAN by Côté et al.
+(2025) or VAE by Jamotton and Hainaut (2024) seem to require significantly more expertise and tuning to get good results.
 Our case study also shows that generic data augmentation does not generally improve the performance of GLMs trained on the real data augmented with the synthetic one.
-We observe that an increase in the proportion of synthetic data compared to the original training data leads to an approximately linear increase in the metrics that measure the difference between coefficients of two GLMs, e.g., the true GLM vs a GLM estimated on augmented data.
-We have noticed only one case where data augmentation improved the estimation of the true GLM coefficients.
+We observe that increasing the proportion of synthetic data relative to the original training data results in approximately linear increases in metrics measuring the difference between the coefficients of two GLMs, e.g., the true GLM vs a GLM estimated on augmented data.
+We have observed only one case in which data augmentation improved the estimation of the true GLM coefficients.
 We leave several aspects for future research.
 For example, it may be interesting to investigate how to integrate business constraints into the amputation-imputation framework based on MICE, e.g., linear constraints on numerical variables or hierarchy among categorical variables.
-Another relevant topic for further research may be the privacy of synthetic datasets generated with the help of MICE.
+Another research direction is the exploration of data augmentation when the original training dataset is small or when the synthetic data generator contains information not present in the training data.
+While ratemaking has historically relied on association measures like correlation, future research must rigorously evaluate the extent to which synthetic data generators preserve the underlying causal structure of insurance risk.
+Finally, future studies should systematically assess the disclosure risk inherent in each generator, specifically quantifying their vulnerability to membership inference and reconstruction attacks.
 Acknowledgments
 Yevhen Havrylenko acknowledges the financial support by the PRIME program (link) of the German Academic Exchange Service (DAAD, https://ror.org/039djdh30) with funds from the German Federal Ministry of Education and Research (BMBF).
-Yevhen Havrylenko acknowledges the support of the University of Copenhagen and Johannes Kepler University Linz, where he conducted parts of this research project.
-Meelis Käärik and Artur Tuttar acknowledge the support of the Estonian Research Council grant PRG1197.
+Yevhen Havrylenko acknowledges the support of Ulm University, the University of Copenhagen and Johannes Kepler University Linz, where he conducted parts of this research project.
+Meelis Käärik and Artur Tuttar acknowledge the support of the Estonian Research Council grants PRG1197 and PRG3105.
 The authors of this publication express their gratitude to Raul Kangro and Märt Möls for their insights, ideas, and discussion related to the article.
-The authors are grateful to the participants of the 28th International Congress on Insurance: Mathematics and Economics.
+The authors are grateful to the participants of the 28th International Congress on Insurance: Mathematics and Economics and 1st ASTIN Bulletin conference.
+The authors are thankful to Magnus Käärik for assistance in refining the language and enhancing clarity.
 Declaration of generative AI and AI-assisted technologies
@@
 The authors declare no conflict of interest.
+Data availability statement
+The authors intend to upload their code and data to a publicly accessible GitHub repository upon acceptance of the article for publication.
 Appendix A Appendix A
@@
   1. 1.
-    use Nearest Centroid Classification to identify for this vector its the closest class; the centroid of each class is determined by the encoding (given by the respective AE fitted before training $GM_{A}$) of the respective category of the original categorical variable;
+    use Nearest Centroid Classification to identify for this vector its closest class; the centroid of each class is determined by the encoding (given by the respective AE fitted before training $GM_{A}$) of the respective category of the original categorical variable;
   2. 2.